### v2

# Frobenius pull-back of parabolic bundles and dormant opers

Abstract.
We study parabolic bundles on an algebraic curve in positive characteristic. Our motivation is to properly formulate Frobenius pull-backs of parabolic bundles in a way that extends various previous facts and arguments for the usual non-parabolic Frobenius pull-backs. After defining that operation, we generalize a classical result by Cartier concerning Frobenius descent, that is, we establish a bijective correspondence (including the version using higher-level $\mathcal{D}$-modules) between parabolic flat bundles with vanishing $p$-curvature on a pointed curve and parabolic bundles on its Frobenius twist. This correspondence gives a description of maximally Frobenius-destabilized parabolic bundles in terms of dormant opers admitting logarithmic poles. As an application of that description together with a previous result in the enumerative geometry of dormant opers, we obtain an explicit formula for computing the number of such parabolic bundles of rank $2$ under certain assumptions.
footnotetext: Y. Wakabayashi: Graduate School of Information Science and Technology, The University of Osaka, Suita, Osaka 565-0871, Japan;footnotetext: e-mail: wakabayashi@ist.osaka-u.ac.jp;footnotetext: 2020 Mathematical Subject Classification: Primary 14H60, Secondary 14G17;footnotetext: Key words: parabolic bundle, positive characteristic, Frobenius pull-back, dormant oper, Frobenius-destabilized bundle
1. Introduction
Let $(g,r)$ be a pair of nonnegative integers with 2g2+r>02g-2+r>0 and $\mathscr{X}:=(X,\{\sigma_{i}\}_{i=1}^{r})$ an $r$-pointed smooth proper curve of genus $g$ over an algebraically closed field $k$. A parabolic (vector) bundle is a vector bundle equipped with weighted flags on its fibers over the marked points $\sigma_{i}$. To be more precise, a parabolic bundle on $\mathscr{X}$ means a collection
$\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ (1.1)
(cf. Definition 2.3) consisting of a vector bundle (i.e., a locally free coherent sheaf) $\mathcal{F}$ on $X$, a set of parabolic weights $\vec{\alpha}$, and an $r$-tuple $(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$, where $\mathfrak{f}_{i}$ ($i=1,\cdots,r$) denotes a sequence of $k$-linear surjections
$\sigma_{i}^{*}\mathcal{F}=:\mathcal{F}_{i}^{[m_{i}]}\twoheadrightarrow\mathcal{F}_{i}^{[m_{i}-1]}\twoheadrightarrow\cdots\twoheadrightarrow\mathcal{F}_{i}^{[1]}\twoheadrightarrow\mathcal{F}_{i}^{[0]}=0$ (1.2)
for some $m_{i}\in\mathbb{Z}_{>0}$. Parabolic bundles and their moduli spaces were introduced by V. B. Mehta and C. S. Seshadri (cf.  [MeSe]), and later investigated by many mathematicians from various points of view, including conformal field theory (cf., e.g.,  [BMW],  [Kum]).
Let us now consider vector bundles in characteristic $p>0$. Note that some remarkable phenomena occurring in characteristic $p$ are derived from properties of the Frobenius morphisms, which are endomorphisms given by sending every regular function to its $p$-th power. For example, a classical result by Cartier (cf.  [Kat, Theorem 5.1]) asserts that the pull-back along the relative Frobenius morphism $F_{X/k}:X\rightarrow X^{(1)}$ from $X$ to its Frobenius twist $X^{(1)}$ over $k$ yields an equivalence of categories
[…]
where a $p$-flat bundle means a vector bundle equipped with (flat) connection with vanishing $p$-curvature (cf.  [Kat, § 5] for the definition of $p$-curvature). This equivalence can be generalized to “higher levels” by using sheaves of differential operators of finite level, introduced by P. Berthelot (cf.  [Ber],  [LeQu, Corollary 3.2.4]). Moreover, we know some extensions of such equivalences, known as the Ogus-Vologodsky correspondence, by which nilpotent Higgs bundles corresponds to flat bundles with nilpotent $p$-curvature (cf.  [GLQ],  [Ohk],  [OgVo],  [Sch]).
In the present paper, we address the issue of extending various results and arguments obtained for the non-parabolic case, including Cartier’s theorem, by appropriately defining “Frobenius pull-back” of parabolic bundles on pointed curves. For positive integers $N$ and $m$, denote by $\Xi_{m,N}^{\leq}$ (cf. (3.11)) the set of $m$-tuples of integers $(a^{[1]},\cdots,a^{[m]})$ with $0\leq a^{[1]}\leq\cdots\leq a^{[m]}<p^{N}$. Let $(m_{1},\cdots,m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an $r$-tuple such that $a_{i}$ (for each $i=1,\cdots,r$) denotes an element $(a_{i}^{[1]},\cdots,a_{i}^{[m_{i}]})$ of $\Xi_{m_{i},N}^{\leq}$. Also, we set $\vec{a}:=(a_{1}/p^{N},\cdots,a_{r}/p^{N})$, where $a_{i}/p^{N}:=(a_{i}^{[1]}/p^{N},\cdots,a_{i}^{[m_{i}]}/p^{N})$. For a parabolic bundle of the form $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ on the $N$-th Frobenius twist $\mathscr{X}^{(N)}:=(X^{(N)},\{\sigma_{i}^{(N)}\}_{i=1}^{r})$ of $\mathscr{X}$, the pull-back of $\mathscr{E}$ along the $N$-th relative Frobenius morphism $F_{X/k}^{(N)}:X\rightarrow X^{(N)}$ yields a $p^{N}$-flat bundle (i.e., a parabolic and level-$N$ enrichment of flat bundle, in the sense of Definition 3.3, (i))
$\mathscr{E}^{F}_{\flat}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{\,F},\vec{a})$ (1.4)
on $\mathscr{X}$ (cf. (4.20)); this will be called the parabolic Frobenius pull-back of $\mathscr{E}$ by $F_{X/k}^{(N)}$ (cf. Definition 4.4). (The pull-backs of parabolic bundles by a ramified covering in characteristic zero are discussed in  [BKP]. Our study can be regarded as its characteristic-$p$ analogue for the case of Frobenius morphisms.) Then, we prove the following assertion generalizing Cartier’s theorem.
Theorem A (cf. Theorem 4.5).
Denote by $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$ (resp., $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}$) the category of parabolic bundles on $\mathscr{X}^{(N)}$ of parabolic weights $\vec{a}/p^{N}$ (resp., parabolic $p^{N}$-flat bundles on $\mathscr{X}$ of parabolic weights $\vec{a}$). Then, the assignment $\mathscr{E}\mapsto\mathscr{E}^{F}_{\flat}$ gives an equivalence of categories
$\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}.$ (1.5)
The second part of the present paper concerns the stability of parabolic bundles. For a parabolic bundle $\mathscr{F}$ as in (1.1), its parabolic slope is defined as the value
$\mathrm{par}\text{-}\mu(\mathscr{F}):=\frac{1}{\mathrm{rk}(\mathcal{F})}\cdot\left(\mathrm{deg}(\mathcal{F})+\sum_{i=1}^{r}\sum_{j}a_{i}^{[j]}\ell_{i}^{[j]}\right),$ (1.6)
where $a_{i}^{[j]}$’s and $\ell_{i}^{[j]}$’s are the parabolic weights and types, respectively, of the parabolic structure $(\vec{\mathfrak{f}},\vec{\alpha})$ at $\sigma_{i}$ (cf. (2.9)). Then, we say that $\mathscr{F}$ is semistable (resp., stable) if the inequality $\mathrm{par}\text{-}\mu(\mathscr{E})\leq\mathrm{par}\text{-}\mu(\mathscr{F})$ (resp., $\mathrm{par}\text{-}\mu(\mathscr{E})<\mathrm{par}\text{-}\mu(\mathscr{F})$) holds for any nontrivial proper parabolic subbundle $\mathscr{E}$ of $\mathscr{F}$ (cf. Definition 2.6).
We are primarily interested in stable parabolic bundles whose parabolic Frobenius pull-back by $F_{X/k}^{(N)}$ are maximally unstable with respect to the Harder-Narasimhan polygon; these bundles are called maximally $F^{(N)}$-destabilized (parabolic) bundles (cf. Definition 6.1). For the previous study of the non-parabolic case, we refer the reader to  [JoPa],  [JRXY],  [Li], and  [Zha]. By a result of K. Joshi and C. Pauly (cf.  [JoPa, Theorem 5.3.1]), the equivalence of categories (1.3) restricts to a bijective correspondence between maximally Frobenius-destabilized bundles and certain $p$-flat bundles called dormant opers (cf.  [JRXY],  [JoPa],  [Jos],  [Moc],  [Wak1],  [Wak2], and  [Wak3] for the previous study of dormant opers).
To state our result concerning its parabolic (and higher-level) generalization, we prepare some notations. Let $n$ be a positive integer, $\mathcal{L}$ a line bundle on $X^{(N)}$, and $\vec{a}$ an element of $(\Xi_{n,N}^{<})^{\times r}$ as in § 5.1 such that $s_{1}^{1}(\vec{a})\in(\Xi_{n,1}^{\leq})^{\times r}$ and $s_{2}^{1}(\vec{a})\in(\Xi_{n,N-1}^{\leq})^{\times r}$ (cf. (4.25) for the definitions of $s_{1}^{1}(-)$ and $s_{2}^{1}(-)$). We set $\vec{1}_{\times n}:=(1_{\times n},\cdots,1_{\times n})$ (= the $r$-tuple of copies of $1_{\times n}:=(1,\cdots,1)\in\Xi_{n,N}^{\leq}$), $\omega:=(n,\vec{1}_{\times n},\vec{a}/p^{N})$, and $\mathcal{L}^{F}_{\flat}:=((F^{(N)*}_{X/k}\mathcal{L})(D_{\star}),\nabla^{\mathrm{can}}_{\mathcal{L},D_{\star}})$ (cf. (3.3)), where $D_{\star}$ denotes the divisor $\sum\limits_{i=1}^{r}(\sum\limits_{j=1}^{n}a_{i}^{[j]})\sigma_{i}$. Then, we obtain the moduli category
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\ \left(\text{resp.,}\ \mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}\right)$ (1.7)
(cf. (5.5), (6.17)) classifying maximally $F^{(N)}$-destablized stable parabolic bundles on $\mathscr{X}^{(N)}$ of rank $n$, type $\vec{1}_{\times n}$, and parabolic weights $\vec{a}/p^{N}$ (resp., dormant $\mathrm{GL}_{n}^{(N)}$-opers on $\mathscr{X}$ of exponent $\vec{a}$ with determinant $\mathcal{L}_{\flat}^{F}$).
Theorem B (cf. Theorem 6.4, (i) and (ii)).
Suppose that the following three conditions are fulfilled:
  • $n\mid\mathrm{deg}(\mathcal{L})$;
  • $\sum\limits_{j=1}^{n}a_{i}^{[j]}<p^{N}$ for every $i=1,\cdots,r$;
  • $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}<\frac{p}{n}$.
Then, there is a natural isomorphism of fibered categories
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\xrightarrow{\sim}\mathcal{O}p_{\mathscr{X},n,\vec{a},\mathcal{L}^{F}_{\flat}}^{{}^{\mathrm{Zzz...}}}$. (1.8)
In particular, there are only finitely many isomorphism classes of maximally $F^{(N)}$-destabilized stable parabolic bundles on $\mathscr{X}^{(N)}$ of rank $n$, type $\vec{1}_{\times n}$, and parabolic weights $\vec{a}/p^{N}$.
As an application of this result and the study on dormant opers discussed in  [Wak3], we can give an explicit formula for computing the number of maximally $F^{(1)}$-destabilized parabolic bundles of rank $2$. The assertion is described as follows.
Theorem C (cf. Theorem 6.7 for the full statement).
Let us take a line bundle $\mathcal{L}$ on $X^{(1)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$. We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$. Suppose further that the following two conditions are fulfilled:
  • $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r<\frac{p}{2}$;
  • $\mathscr{X}$ is sufficiently general in the moduli stack $\mathcal{M}_{g,r}$ of $r$-pointed smooth proper curves of genus $g$.
Then, the cardinality $\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))$ of the set of isomorphism classes of parabolic bundles classified by $\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k)$ satisfies the following equality:
$\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))=2\cdot p^{g-1}\cdot\sum_{j=1}^{p-1}\frac{\prod\limits_{i=1}^{r}(-1)^{(j+1)(a_{i}^{[2]}-a_{i}^{[1]}+1)}\sin\left(\frac{(a_{i}^{[2]}-a_{i}^{[1]})j\pi}{p}\right)}{\sin^{2g-2+r}\left(\frac{j\pi}{p}\right)}.$ (1.9)
Finally, we remark that parabolic bundles in positive characteristic were previously investigated in  [KuMa] and  [KuPa] (resp.,  [LiSu] and  [She]) from a viewpoint of Galois coverings (resp., non-Abelian Hodge correspondence). It should be a future task to understand how can we relate that work to our discussion concerning parabolic Frobenius pull-back.
Notation and Conventions
Throughout the present paper, we fix a prime number $p$, an algebraically closed field $k$ of characteristic $p$, and a pair of nonnegative integers $(g,r)$. Also, we fix a $k$-scheme $S$ and an $r$-pointed smooth curve
$\mathscr{X}:=(f:X\rightarrow S,\{\sigma_{i}:S\rightarrow X\}_{i=1}^{r})$ (1.10)
of genus $g$ over $S$, i.e., $X$ is a smooth proper curve over $S$ of genus $g$ and $\{\sigma_{i}\}_{i=1}^{r}$ denotes a collection of mutually disjoint $r$ marked points. If $S^{\prime}$ is an $S$-scheme, then we denote by $\mathscr{X}_{S^{\prime}}$ the base-change of $\mathscr{X}$ over $S^{\prime}$.
Denote by $D$ the effective relative divisor on $X$ (relative to $S$) defined as the union of the marked points $\{\sigma_{i}\}_{i=1}^{r}$; it defines a log structure on $X$, and the resulting log scheme will be denoted by $X^{\mathrm{log}}$.
Since $X^{\mathrm{log}}/S$ is log smooth, the sheaf of logarithmic $1$-forms $\Omega_{X^{\mathrm{log}}/S}$ of $X^{\mathrm{log}}/S$, as well as its dual $\mathcal{T}_{X^{\mathrm{log}}/S}:=\Omega_{X^{\mathrm{log}}/S}^{\vee}$, is a line bundle. In fact, we have natural isomorphisms $\Omega_{X^{\mathrm{log}}/S}\cong\Omega_{X/S}(D)$, $\mathcal{T}_{X^{\mathrm{log}}/S}\cong\mathcal{T}_{X/S}(-D)$. When there is no fear of confusion, we write $\Omega$ and $\mathcal{T}$ instead of $\Omega_{X^{\mathrm{log}}/S}$ and $\mathcal{T}_{X^{\mathrm{log}}/S}$, respectively. Also, for a positive integer $N$, we denote by $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$ (cf.  [Mon, Definition 2.3.1]) the sheaf of logarithmic differential operators on $X^{\mathrm{log}}/S$ of level N1N-1.
Next, write $F_{S}$ (resp., $F_{X}$) for the absolute Frobenius endomorphism of $S$ (resp., $X$). For each positive integer $N$, we shall denote by $X^{(N)}$ the base-change $S\times_{F_{S}^{N},S}X$ of $X$ along the $N$-th iterate $F_{S}^{N}$ of $F_{S}$; we will refer to it as the $N$-th Frobenius twist of $X$ over $S$. Also, the morphism $F_{X/S}^{(N)}\left(:=(f,F_{X}^{N})\right):X\rightarrow X^{(N)}$ is called the $N$-th relative Frobenius morphism of $X$ over $S$.
2. Parabolic vector bundles
In this section, we recall parabolic bundles on an algebraic curve and the stability for such objects. There are many references discussing parabolic bundles; for example, we refer the reader to  [MaYo],  [MeSe].
2.1. Definition of a parabolic bundle
In Definitions 2.1 and 2.2 below, we shall fix a vector bundle $\mathcal{F}$ on $X$ of constant rank.
Definition 2.1.
  • (i)
    Let $i$ be an element of $\{1,\cdots,r\}$, and we shall write $\mathcal{F}_{i}:=\sigma_{i}^{*}\mathcal{F}$, which forms a vector bundle on $S$. By a quasi-parabolic structure of $\mathcal{F}$ at $\sigma_{i}$, we mean a sequence of $\mathcal{O}_{S}$-linear surjections
    $\mathfrak{f}_{i}:\mathcal{F}_{i}=\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\mathfrak{f}_{i}^{[m_{i}]}}\mathcal{F}_{i}^{[m_{i}-1]}\xrightarrow{\mathfrak{f}_{i}^{[m_{i}-1]}}\cdots\xrightarrow{\mathfrak{f}_{i}^{[2]}}\mathcal{F}_{i}^{[1]}\xrightarrow{\mathfrak{f}_{i}^{[1]}}\mathcal{F}_{i}^{[0]}=0$ (2.1)
    (where $m_{i}\in\mathbb{Z}_{>0}$) between vector bundles $\mathcal{F}_{i}^{[j]}$ on $S$. A parabolic structure of $\mathcal{F}$ at $\sigma_{i}$ is a pair
    $(\mathfrak{f}_{i},\alpha_{i})$ (2.2)
[…]
  • (ii)
    A quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ is a collection of data
    $\vec{\mathfrak{f}}\ \left(\text{resp.,}\ (\vec{\mathfrak{f}},\vec{\alpha})\right)$ (2.3)
    consisting of an $r$-tuple $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ (resp., $r$-tuples $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$) such that, for each $i=1,\cdots,r$, $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$) forms a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$.
Definition 2.2.
  • (i)
    Let $i$ be an element of $\{1,\cdots,r\}$ and $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$) a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$ at $\sigma_{i}$ as in (2.2). Then, the integer $m_{i}$ is called the length of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$). Also, if $S$ is connected, then we obtain the $m_{i}$-tuple of nonnegative integers
    $\ell_{i}:=(\ell_{i}^{[1]},\cdots,\ell_{i}^{[m_{i}]}),$ (2.4)
    where $\ell_{i}^{[j]}$ denotes the rank of the vector bundle $\mathrm{Ker}(\mathfrak{f}_{i}^{[j]})$, and it is called the type of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$).
  • (ii)
    Let $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$), where $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ (resp., $\vec{\mathfrak{f}}:=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$ and $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$), be a quasi-parabolic (resp., parabolic) structure of $\mathcal{F}$. Under the assumption that $S$ is connected, we denote by $\ell_{i}$ the type of $\mathfrak{f}_{i}$ (resp., $(\mathfrak{f}_{i},\alpha_{i})$). Then, the collections
    $\vec{m}:=(m_{1},\cdots,m_{r})\ \text{and}\ \vec{\ell}:=(\ell_{1},\cdots,\ell_{r})$ (2.5)
    are called the length and type of $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$), respectively.
Definition 2.3.
  • (i)
    A quasi-parabolic (resp., parabolic) bundle on $\mathscr{X}$ is a collection of data
    $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})\ \left(\text{resp.,}\ \mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})\right),$ (2.6)
    consisting of a vector bundle $\mathcal{F}$ on $X$ of constant rank and a quasi-parabolic (resp., parabolic) structure $\vec{\mathfrak{f}}$ (resp., $(\vec{\mathfrak{f}},\vec{\alpha})$) of $\mathcal{F}$. If $\mathfrak{f}_{i}:=\{\mathfrak{f}_{i}^{[j^{\prime}]}\}_{j^{\prime}=1}^{m_{i}}$ denotes the $i$-th factor of $\vec{\mathfrak{f}}$ (where $i\in\{1,\cdots,r\}$), then, for each $j=1,\cdots,m_{i}$, we obtain a composite surjection
    $\varpi_{\mathscr{F},i}^{[j]}:\mathcal{F}\twoheadrightarrow\sigma_{i*}\mathcal{F}_{i}^{[m_{i}]}\xrightarrow{\sigma_{i*}\mathfrak{f}_{i}^{[m_{i}]}}\cdots\xrightarrow{\sigma_{i*}\mathfrak{f}_{i}^{[j+1]}}\sigma_{i*}\mathcal{F}_{i}^{[j]}.$ (2.7)
  • (ii)
    Let $\mathscr{F}$ be a (quasi-)parabolic bundle on $\mathscr{X}$ as in (2.6). We shall say that $\mathscr{F}$ is of rank $n(>0)$ if the vector bundle $\mathcal{F}$ has constant rank $n$. Also, in the resp’d portion of (i), the numbers in $\vec{\alpha}$ are called the parabolic weights of $\mathscr{F}$.
Remark 2.4.
For convenience of our discussion (cf. Definition 3.3, (i)), we do not impose (unlike the usual definition) the condition that the parabolic weights of a parabolic bundle is less than $1$.
Next, let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{\alpha})$ be two parabolic bundles on $\mathscr{X}$ with parabolic structures having common length and parabolic weights
Definition 2.5.
A morphism of parabolic bundles from $\mathscr{E}$ to $\mathscr{F}$ is an $\mathcal{O}_{X}$-linear morphism $h:\mathcal{E}\rightarrow\mathcal{F}$ that restricts to a morphism $\mathrm{Ker}(\varpi_{\mathscr{E},i}^{[j]})\rightarrow\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ for every $i$ and $j$.
Let us choose an $r$-tuple $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$ such that $\alpha_{i}$ (for each $i=1,\cdots,r$) is a collection of real numbers $(\alpha_{i}^{[1]},\cdots,\alpha_{i}^{[m_{i}]})$ ($m_{i}\in\mathbb{Z}_{>0}$) with $0\leq\alpha_{i}^{[1]}\leq\cdots\leq\alpha_{i}^{[m_{i}]}$. This $r$-tuple determines the category
$\mathcal{B}un_{\mathscr{X},\vec{\alpha}}$ (2.8)
consisting of parabolic bundles on $\mathscr{X}$ having (length $(m_{1},\cdots,m_{r})$ and) parabolic weights $\vec{\alpha}$.
2.2. Stability for parabolic bundles
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha})$ be a parabolic bundle on $\mathscr{X}$ such that $(\vec{\mathfrak{f}},\vec{\alpha})$ is a collection as in the resp’d portion of (2.3). Given a subbundle $\mathcal{E}$ of $\mathcal{F}$ of constant rank, we set $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ (for each $i=1,\cdots,r$ and $j=1,\cdots,m_{i}$) to be the image of $\mathcal{E}$ via $\varpi_{\mathscr{F},i}^{[j]}$. Assume that $\mathcal{F}^{[j]}_{i,\mathcal{E}}$ are all vector bundles on $S$ when restricted via $\sigma_{i}$. (Of course, this assumption is automatically satisfied if $S=\mathrm{Spec}(k)$.) Since $\mathfrak{f}_{i}^{[j]}$ restricts to a surjection $\mathfrak{f}_{i,\mathcal{E}}^{[j]}:\mathcal{F}^{[j]}_{i,\mathcal{E}}\twoheadrightarrow\mathcal{F}^{[j-1]}_{i,\mathcal{E}}$, the set of morphisms $\mathfrak{f}_{i,\mathcal{E}}:=\{\mathfrak{f}_{i,\mathcal{E}}^{[j]}\}_{j=1}^{m_{i}}$ specifies, via $\sigma_{i}^{*}(-)$, a quasi-parabolic structure on $\mathcal{E}$ at $\sigma_{i}$. The resulting collection $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{\alpha})$, where $\vec{\mathfrak{f}}_{\mathcal{E}}:=(\mathfrak{f}_{1,\mathcal{E}},\cdots,\mathfrak{f}_{r,\mathcal{E}})$, forms a parabolic bundle on $\mathscr{X}$; such a parabolic bundle obtained in this way is called parabolic subbundle of $\mathscr{F}$.
Suppose that $S$ is connected. Each parabolic subbundle $\mathscr{E}$ as above determines the following two real numbers:
$\mathrm{par}\text{-}\mathrm{deg}(\mathscr{E})=\mathrm{deg}(\mathcal{E})+\sum_{i=1}^{r}\sum_{j=1}^{m_{i}}\alpha_{i}^{[j]}\ell_{i,\mathcal{E}}^{[j]},\hskip 22.76219pt\mathrm{par}\text{-}\mu(\mathscr{E}):=\frac{\mathrm{par}\text{-}\mathrm{deg}(\mathscr{E})}{\mathrm{rk}(\mathcal{E})},$ (2.9)
[…]
Definition 2.6.
We shall say that $\mathscr{F}$ is semistable (resp., stable) if for any geometric point $t$ of $S$ and any nontrivial proper parabolic subbundle $\mathscr{E}_{t}$ of the fiber $\mathscr{F}_{t}$ of $\mathscr{F}$ over $t$, one has
$\mathrm{par}\text{-}\mu(\mathscr{E}_{t})\leq\mathrm{par}\text{-}\mu(\mathscr{F}_{t})\ \left(\text{resp.,}\ \mathrm{par}\text{-}\mu(\mathscr{E}_{t})<\mathrm{par}\text{-}\mu(\mathscr{F}_{t})\right).$ (2.10)
Let $\omega:=(n,\vec{\ell},\vec{\alpha})$ be a triple consisting of a positive integer $n$, an $r$-tuple $\vec{\ell}:=(\ell_{1},\cdots,\ell_{r})$ such that $\ell_{i}$ (for each $i=1,\cdots,r$) is a sequence of nonnegative integers $(\ell_{i}^{[1]},\cdots,\ell_{i}^{[m_{i}]})$ (where $m_{i}\in\mathbb{Z}_{>0}$), and a collection of parabolic weights $\vec{\alpha}:=(\alpha_{1},\cdots,\alpha_{r})$ as before. Then, for a line bundle $\mathcal{L}$ on $X$, we obtain the category fibered in groupoids
$\overline{\mathcal{U}}_{\mathscr{X},\omega,\mathcal{L}}\ \left(\text{resp.,}\ \mathcal{U}_{\mathscr{X},\omega,\mathcal{L}}\right)$ (2.11)
over the category of $S$-schemes $\mathcal{S}ch_{/S}$ defined as follows: the fiber $\overline{\mathcal{U}}_{\mathscr{X},\omega,\mathcal{L}}(S^{\prime})$ (resp., $\mathcal{U}_{\mathscr{X},\omega,\mathcal{L}}(S^{\prime})$) over each $S$-scheme $s:S^{\prime}\rightarrow S$ is the groupoid of pairs $(\mathscr{F},\eta)$, where
  • $\mathscr{F}$ denotes a rank-$n$ semistable (resp., stable) parabolic bundle on $\mathscr{X}_{S^{\prime}}$ whose parabolic structure has type $\vec{\ell}$ and parabolic weights $\vec{\alpha}$;
  • $\eta$ denotes an isomorphism $\mathrm{det}(\mathcal{F})\xrightarrow{\sim}\mathcal{L}$, where $\mathcal{F}$ is the underlying vector bundle of $\mathscr{F}$.
3. Parabolic flat bundles of higher level
This section deals with a higher-level generalization of $\mathcal{D}$-modules in positive characteristic. This concept was originally studied in  [Ber] and later generalized in  [Mon] to the logarithmic situation. We here recall monodromy operators (in other words, residues) and exponents of higher level, in the manner of  [Wak4]; the exponents of a higher-level $\mathcal{D}$-module describe its local structures around the marked points.
3.1. Log flat bundles and higher-level $\mathcal{D}$-modules
Let us fix a positive integer $N$. For simplicity, we use the notation $\mathcal{D}^{(N-1)}$” to denote the sheaf $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}/S}$. For each integer $j$, we shall write $\mathcal{D}^{(N-1)}_{<j}$ for the subsheaf of $\mathcal{D}^{(N-1)}$ consisting of logarithmic differential operators of order <j<j.
Note that $\mathcal{D}^{(N-1)}$ is endowed with a structure of $\mathcal{O}_{X}$-module arising from left (resp., right) multiplication by sections of $\mathcal{D}_{<1}^{(N-1)}\left(=\mathcal{O}_{X}\right)$; we denote the resulting $\mathcal{O}_{X}$-module by ${{}^{L}}\mathcal{D}^{(N-1)}$ (resp., ${{}^{R}}\mathcal{D}^{(N-1)}$). A (left) $\mathcal{D}^{(N-1)}$-module structure on an $\mathcal{O}_{X}$-module $\mathcal{F}$ is a left $\mathcal{D}^{(N-1)}$-action $\nabla:{{}^{L}}\mathcal{D}^{(N-1)}\rightarrow\mathcal{E}nd_{f^{-1}(\mathcal{O}_{S})}(\mathcal{F})$ on $\mathcal{F}$ extending its $\mathcal{O}_{X}$-module structure. An $\mathcal{O}_{X}$-module equipped with a $\mathcal{D}^{(N-1)}$-module structure is called a (left) $\mathcal{D}^{(N-1)}$-module.
Given a $\mathcal{D}^{(N-1)}$-module $(\mathcal{F},\nabla)$, we shall write $\mathcal{S}ol(\nabla)$ for the subsheaf of $\mathcal{F}$ on which $\mathcal{D}_{+}^{(N-1)}$ acts as zero, where $\mathcal{D}_{+}^{(N-1)}$ denotes the kernel of the canonical projection $\mathcal{D}^{(N-1)}\twoheadrightarrow\mathcal{O}_{X}$. The sheaf $\mathcal{S}ol(\nabla)$ may be regarded as an $\mathcal{O}_{X^{(N)}}$-module via the underlying homeomorphism of $F_{X/S}^{(N)}$.
Denote by $\psi$ the $p^{N}$-curvature map $\mathcal{T}^{\otimes p^{N}}\rightarrow\mathcal{D}^{(N-1)}$ defined in  [Ohk, § 3.2, Definition 3.10]. Given a $\mathcal{D}^{(N-1)}$-module $(\mathcal{F},\nabla)$, we shall set
$\psi_{(\mathcal{F},\nabla)}:=\nabla\circ\psi:\mathcal{T}^{\otimes p^{N}}\rightarrow\mathcal{E}nd_{f^{-1}(\mathcal{O}_{S})}(\mathcal{F}),$ (3.1)
which will be called the $p^{N}$-curvature of $(\mathcal{F},\nabla)$. It is well-known that the image of $\psi_{(\mathcal{F},\nabla)}$ lies in $\mathcal{E}nd_{\mathcal{O}_{X}}(\mathcal{F})$.
Definition 3.1.
By a $p^{N}$-flat bundle on $\mathscr{X}$, we shall mean a pair $(\mathcal{F},\nabla)$ consisting of a vector bundle $\mathcal{F}$ on $X$ and a $\mathcal{D}^{(N-1)}$-module structure $\nabla$ on $\mathcal{F}$ with vanishing $p^{N}$-curvature.
Let $\mathcal{E}$ be an $\mathcal{O}_{X^{(N)}}$-module. According to  [Mon, Corollaire 3.3.1], there exists a canonical $\mathcal{D}^{(N-1)}$-module structure
$\nabla_{\mathcal{E}}^{\mathrm{can}}:{{}^{L}}\mathcal{D}^{(N-1)}\rightarrow\mathcal{E}nd_{f^{-1}(\mathcal{O}_{S})}(F^{(N)*}_{X/S}\mathcal{E})$ (3.2)
on the $N$-th Frobenius pull-back $F^{(N)*}_{X/S}\mathcal{E}$ uniquely determined by the condition that every local section $(F^{(N)}_{X/S})^{-1}(v)$ (for $v\in\mathcal{E}$) is horizontal. Moreover, for a relative divisor $D_{0}$ on $X/S$ with $\mathrm{Supp}(D_{0})\subseteq\mathrm{Supp}(D)\left(=\bigcup_{i=1}^{r}\mathrm{Im}(\sigma_{i})\right)$, one can obtain a $\mathcal{D}^{(N-1)}$-module structure
$\nabla_{\mathcal{E},D_{0}}^{\mathrm{can}}:{{}^{L}}\mathcal{D}^{(N-1)}\rightarrow\mathcal{E}nd_{f^{-1}(\mathcal{O}_{S})}((F^{(N)*}_{X/S}\mathcal{E})(D_{0}))$ (3.3)
on $(F^{(N)*}_{X/S}\mathcal{E})(D_{0})\left(=F^{(N)*}_{X/S}\mathcal{E}\otimes\mathcal{O}_{X}(D_{0})\right)$ that is consistent with $\nabla_{\mathcal{E}}^{\mathrm{can}}$ outside $\mathrm{Supp}(D_{0})$. It is immediately verified that $\nabla^{\mathrm{can}}_{\mathcal{E}}$ and $\nabla_{\mathcal{E},D_{0}}^{\mathrm{can}}$ have vanishing $p^{N}$-curvature.
3.2. Monodromy operators
We shall set
$\mathcal{B}_{S}:=\bigoplus_{j\in\mathbb{Z}_{\geq 0}}\mathcal{O}_{S}\cdot\partial_{\mathcal{B}}^{\langle j\rangle},$ (3.4)
where $\partial_{\mathcal{B}}^{\langle i\rangle}$’s are abstract symbols (cf.  [Wak4, § 4.2]). Here, for each $j\in\mathbb{Z}_{\geq 0}$, let $q_{j}$ denote a unique nonnegative integer satisfying $j=p^{N-1}\cdot q_{j}+r_{j}$ and $0\leq r_{j}<p^{N-1}$. Then, we equip $\mathcal{B}_{S}$ with a structure of $\mathcal{O}_{S}$-algebra given by
$\partial_{\mathcal{B}}^{\langle j^{\prime}\rangle}\cdot\partial_{\mathcal{B}}^{\langle j^{\prime\prime}\rangle}:=\sum_{j=\mathrm{max}\{j^{\prime},j^{\prime\prime}\}}^{j^{\prime}+j^{\prime\prime}}\frac{j!}{(j^{\prime}+j^{\prime\prime}-j)!\cdot(j-j^{\prime})!\cdot(j-j^{\prime\prime})!}\cdot\frac{q_{j^{\prime}}!\cdot q_{j^{\prime\prime}}!}{q_{j}!}\cdot\partial_{\mathcal{B}}^{\langle j\rangle}.$ (3.5)
The $\mathcal{O}_{S}$-algebra $\mathcal{B}_{S}$ is commutative and generated by the global sections $\partial_{\mathcal{B}}^{\langle 1\rangle},\partial_{\mathcal{B}}^{\langle p\rangle},\cdots,\partial_{\mathcal{B}}^{\langle p^{N-1}\rangle}$.
Let $\mathcal{G}$ be an $\mathcal{O}_{S}$-module and $\mu$ a morphism of $\mathcal{O}_{S}$-algebras $\mathcal{B}_{S}\rightarrow\mathcal{E}nd_{\mathcal{O}_{S}}(\mathcal{G})$. For each $a\in\mathbb{Z}_{\geq 0}$, we set $\mu^{\langle p^{a}\rangle}:=\mu(\partial_{\mathcal{B}}^{\langle p^{a}\rangle})$, which is an element of $\mathrm{End}_{\mathcal{O}_{S}}(\mathcal{G})$. The morphism $\mu$ is uniquely determined by the $N$-tuple
$\mu^{\langle\bullet\rangle}:=(\mu^{\langle 1\rangle},\mu^{\langle p\rangle}\cdots,\mu^{\langle p^{N-1}\rangle})\in\mathrm{End}_{\mathcal{O}_{S}}(\mathcal{G})^{\oplus N}.$ (3.6)
[…]
$\mu_{i}(\nabla):\mathcal{B}_{S}\xrightarrow{\sim}\sigma_{i}^{*}({{}^{L}}\mathcal{D}^{(N-1)})\xrightarrow{\sigma_{i}^{*}\nabla}\mathcal{E}nd_{\mathcal{O}_{S}}(\sigma_{i}^{*}\mathcal{F}),$ (3.7)
and the pair $(\sigma_{i}^{*}\mathcal{F},\mu_{i}(\nabla))$ specifies a $\mathcal{B}_{S}$-module. In particular, we obtain
$\mu_{i}(\nabla)^{\langle\bullet\rangle}:=(\mu_{i}(\nabla)^{\langle 1\rangle},\mu_{i}(\nabla)^{\langle p\rangle}\cdots,\mu_{i}(\nabla)^{\langle p^{N-1}\rangle})\in\mathrm{End}_{\mathcal{O}_{S}}(\sigma_{i}^{*}\mathcal{F})^{\oplus N}.$ (3.8)
Since $\mathcal{B}_{S}$ is commutative, the elements $\mu_{i}(\nabla)^{\langle 1\rangle},\cdots,\mu_{i}(\nabla)^{\langle p^{N-1}\rangle}$ commute with each other. We refer to $\mu_{i}(\nabla)$ and $\mu_{i}(\nabla)^{\langle\bullet\rangle}$ as the monodromy operator of $\nabla$ at $\sigma_{i}$. In the case where $\mathcal{F}$ is a line bundle, $\mu_{i}(\nabla)^{\langle\bullet\rangle}$ may be regarded as an element of $\Gamma(S,\mathcal{O}_{S})^{\oplus N}\left(=\mathrm{End}_{\mathcal{O}_{S}}(\sigma_{i}^{*}\mathcal{F})^{\oplus N}\right)$.
3.3. Parabolic $p^{N}$-flat bundles
In what follows, we generalize the classical definition of a parabolic connection to our situation.
Definition 3.2.
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$. A $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$ is a $\mathcal{D}^{(N-1)}$-module structure $\nabla$ on $\mathcal{F}$ such that $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$ for every $i=1,\cdots,r$ and $j=0,\cdots,m_{i}$.
Let $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}})$ be a quasi-parabolic bundle on $\mathscr{X}$ and $(m_{1},\cdots,m_{r})$ the length of $\vec{\mathfrak{f}}=(\mathfrak{f}_{1},\cdots,\mathfrak{f}_{r})$. Also, let $\nabla$ be a $\mathcal{D}^{(N-1)}$-module structure on $\mathscr{F}$. Since $\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})$ (for each $i=1,\cdots,r$ and $j=1,\cdots m_{i}$) is closed under $\nabla$, we see that $\nabla$ induces a $\mathcal{B}_{S}$-module structure
$\mu_{i}^{[j]}(\nabla):\mathcal{B}_{S}\rightarrow\mathcal{E}nd_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))$ (3.9)
on $\mathrm{Ker}(\mathfrak{f}_{i}^{[j]})\left(\cong\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j-1]})/\mathrm{Ker}(\varpi_{\mathscr{F},i}^{[j]})\right)$; this corresponds to
$\mu_{i}^{[j]}(\nabla)^{\langle\bullet\rangle}:=(\mu_{i}^{[j]}(\nabla)^{\langle 1\rangle},\mu_{i}^{[j]}(\nabla)^{\langle p\rangle},\cdots,\mu_{i}^{[j]}(\nabla)^{\langle p^{N-1}\rangle})\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))^{\oplus N}.$ (3.10)
Here, for each $m\in\mathbb{Z}_{>0}$, we shall denote by
$\Xi_{m,N}^{\leq}\ \left(\text{resp.,}\ \Xi_{m,N}^{<}\right)$ (3.11)
the set of $m$-tuples of integers $(a^{[1]},\cdots,a^{[m]})$ with $0\leq a^{[1]}\leq\cdots\leq a^{[m]}<p^{N}$ (resp., $0\leq a^{[1]}<\cdots<a^{[m]}<p^{N}$). Also, given an element $d$ of $\mathbb{Z}/p^{N}\mathbb{Z}$, we denote by $\widetilde{d}$ the integer defined as the unique lifting of $d$ via the natural surjection $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$ satisfying $0\leq\widetilde{d}<p^{N}$. Let $\widetilde{d}_{[0]},\cdots,\widetilde{d}_{[N-1]}$ be the collection of integers uniquely determined by the condition that $\widetilde{d}=\sum\limits_{s=0}^{N-1}p^{s}\cdot\widetilde{d}_{[s]}$ and $0\leq\widetilde{d}_{[s]}<p$ ($s=0,\cdots,N-1$). For each $s=0,\cdots,N-1$, we write $\widetilde{d}_{[0,s]}:=\sum\limits_{s^{\prime}=0}^{s}p^{s^{\prime}}\cdot\widetilde{d}_{[s^{\prime}]}$ and write $d_{[s]}$ (resp., $d_{[0,s]}$) for the image of $\widetilde{d}_{[s]}$ (resp., $\widetilde{d}_{[0,s]}$) via the natural projection $\mathbb{Z}\twoheadrightarrow\mathbb{F}_{p}$ (resp., $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{s+1}\mathbb{Z}$).
Definition 3.3.
  • (i)
    Let us consider a collection of data
    $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ (3.12)
    where
    • $(\mathcal{F},\vec{\mathfrak{f}},\vec{a})$ denotes a parabolic bundle on $\mathscr{X}$ such that $\vec{a}$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ for some $(m_{1},\cdots,m_{r})\in\mathbb{Z}_{>0}^{\times r}$ (in particular, $(m_{1},\cdots,m_{r})$ coincides with the length of $\vec{\mathfrak{f}}$);
    • $\nabla$ denotes a $\mathcal{D}^{(N-1)}$-module structure on $(\mathcal{F},\vec{\mathfrak{f}})$ with $\psi_{(\mathcal{F},\nabla)}=0$ (resp., $(\psi_{(\mathcal{F},\nabla)})^{p-1}=0$).
    Then, we say that $\mathscr{F}_{\flat}$ is a parabolic $p^{N}$-flat (resp., parabolic $p^{N}$-nilpotent) bundle on $\mathscr{X}$ if, for any $i=1,\cdots,r$, $j=1,\cdots,m_{i}$, and $s=0,\cdots,N-1$, the element $\mu_{i}^{[j]}(\nabla)^{\langle p^{s}\rangle}\in\mathrm{End}_{\mathcal{O}_{S}}(\mathrm{Ker}(\mathfrak{f}_{i}^{[j]}))$ coincides with the multiplication by $(-a_{i}^{[j]})_{[s]}\in\mathbb{F}_{p}$.
    The notion of a morphism between parabolic $p^{N}$-flat (resp., parabolic $p^{N}$-nilpotent) bundles can be formulated naturally.
  • (iii)
    Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ be a parabolic $p^{N}$-flat bundle on $\mathscr{X}$. By a parabolic $p^{N}$-flat subbundle of $\mathscr{F}_{\flat}$, we mean a parabolic $p^{N}$-flat bundle of the form $\mathscr{E}_{\flat}:=(\mathcal{E},\nabla|_{\mathcal{E}},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{a})$, where $(\mathcal{E},\vec{\mathfrak{f}}_{\mathcal{E}},\vec{a})$ is a parabolic subbundle of $\mathscr{F}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{a})$ such that $\mathcal{E}$ is closed under the $\mathcal{D}^{(N-1)}$-action $\nabla$, and $\nabla|_{\mathcal{E}}$ denotes the resulting $\mathcal{D}^{(N-1)}$-action induced by $\nabla$.
Each element $\vec{a}:=(a_{1},\cdots,a_{r})$ of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ (where $m_{1},\cdots,m_{r}\in\mathbb{Z}_{>0}$) determines the category
$\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un^{\psi=0}_{\mathscr{X},\vec{a}}\ \left(\text{resp.,}\ \mathcal{D}^{(N-1)}\text{-}\mathcal{B}un^{\mathrm{nilp}}_{\mathscr{X},\vec{a}}\right)$ (3.13)
of parabolic $p^{N}$-flat (resp., $p^{N}$-nilpotent) bundles on $\mathscr{X}$ of (length $(m_{1},\cdots,m_{r})$ and) parabolic weights $\vec{a}$.
3.4. Exponent of a $p^{N}$-flat bundle
Consider the formal disc
$U:=\mathcal{S}pec(\mathcal{O}_{S}[\![t]\!])$ (3.14)
over $S$, where $t$ denotes a formal parameter; it has a closed immersion $\sigma_{\circ}:S\hookrightarrow U$ determined by “t=0t=0”. The pair
$\mathscr{U}:=(U,\sigma_{\circ})$ (3.15)
may be regarded as a “local curve” over $S$. In fact, the formal completion of $\mathscr{X}$ along the marked point is, locally on $S$, isomorphic to $U$.
Note that various formulations and definitions discussed so far are available when $\mathscr{X}$ is replaced with $\mathscr{U}$. In particular, if $U^{\mathrm{log}}$ denotes the log scheme $U$ equipped with the log structure given by the relative divisor $D_{\circ}:=\mathrm{Im}(\sigma_{\circ})$, then we obtain $\mathcal{D}^{(N-1)}_{\circ}:=\mathcal{D}^{(N-1)}_{U^{\mathrm{log}}/S}$. The $\mathcal{O}_{U}$-module ${{}^{L}}\mathcal{D}_{\circ}^{(m)}$ decomposes into the direct sum $\bigoplus_{j\in\mathbb{Z}_{\geq 0}}\mathcal{O}_{U}\cdot\partial^{\langle j\rangle}_{\circ}$, where $\{\partial^{\langle j\rangle}_{\circ}\}_{j}$ is the basis associated to the logarithmic coordinate determined by $t$.
For $d\in\mathbb{Z}/p^{N}\mathbb{Z}$, there exists a unique $\mathcal{D}_{\circ}^{(N-1)}$-module structure
$\nabla_{d}:\mathcal{D}_{\circ}^{(N-1)}\rightarrow\mathcal{E}nd_{\mathcal{O}_{S}}(\mathcal{O}_{U})$ (3.16)
on $\mathcal{O}_{U}$ determined by the condition that $\nabla_{d}(\partial^{\langle j\rangle}_{\circ})(t^{n})=q_{j}!\cdot\binom{n-\widetilde{d}}{j}\cdot t^{n}$ for every $j$, $n\in\mathbb{Z}_{\geq 0}$. The resulting $\mathcal{D}_{\circ}^{(N-1)}$-module
$\mathscr{O}_{d,\flat}:=(\mathcal{O}_{U},\nabla_{d})$ (3.17)
is isomorphic to the unique extension of $\mathscr{O}_{0,\flat}$ to $t^{-\widetilde{d}}\cdot\mathcal{O}_{U}\left(\supseteq\mathcal{O}_{U}\right)$. In particular, $\nabla_{d}$ has vanishing $p^{N}$-curvature. Also, it follows from  [Wak4, Proposition 4.3.2] that the equality
$\mu(\nabla_{d})^{\langle\bullet\rangle}=((-d)_{[0]},\cdots,(-d)_{[N-1]})$ (3.18)
holds under the natural identification $\mathrm{End}_{\mathcal{O}_{S}}(\sigma_{\circ}^{*}\mathcal{O}_{U})=\Gamma(S,\mathcal{O}_{S})\left(\supseteq\mathbb{F}_{p}\right)$.
Let $(\mathcal{F},\nabla)$ be a $\mathcal{D}_{\circ}^{(N-1)}$-module with vanishing $p^{N}$-curvature such that $\mathcal{F}$ is a vector bundle of rank $n>0$. Suppose first that $S=\mathrm{Spec}(R)$ for a local ring $(R,\mathfrak{m})$ over $k$ with $R/\mathfrak{m}=k$. According to  [Wak4, Proposition-Definition 4.4.1], there exists an isomorphism of $\mathcal{D}_{\circ}^{(N-1)}$-modules
$\bigoplus_{j=1}^{n}\mathscr{O}_{d^{[j]},\flat}\xrightarrow{\sim}(\mathcal{F},\nabla)$ (3.19)
for some $d^{[1]},\cdots,d^{[n]}\in\mathbb{Z}/p^{N}\mathbb{Z}$; the multiset $d^{[1]},\cdots,d^{[n]}]$ depends only on the isomorphism class of $(\mathcal{F},\nabla)$. Even in the case where $S$ is an arbitrary connected $k$-scheme (that is not necessarily the spectrum of a local ring), we can find a well-defined multiset of integers
$e(\nabla):=[a^{[1]},\cdots,a^{[n]}]$ (3.20)
in $\{0,1,\cdots,p^{N}-1\}$ satisfying the following condition: if $d^{[j]}$ denotes the image of $a^{[j]}$ via $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$ (hence $\widetilde{d}^{[j]}=a^{[j]}$), then $(\mathcal{F},\nabla)$ admits an isomorphism as in (3.19) when restricted over the spectrum of the local ring determined by every closed point of $S$.
Definition 3.4.
We refer to $e(\nabla)$ as the exponent of $(\mathcal{F},\nabla)$ (or, of $\nabla$). (This definition is equivalent to the one discussed in  [Wak4, Proposition-Definition 4.4.1] under the identification $d\leftrightarrow\widetilde{d}$.)
Definition 3.5.
Let $(\mathcal{F},\nabla)$ be a $p^{N}$-flat bundle on $\mathscr{X}$ and $i$ an element of $\{1,\cdots,r\}$. Also, let $e$ be a multiset of elements in $\{0,\cdots,p^{N}-1\}$ with cardinality $n$. Then, we say that $(\mathcal{F},\nabla)$ has exponent $e$ at $\sigma_{i}$ if, locally on $S$, the restriction of $(\mathcal{F},\nabla)$ to the formal completion $\mathscr{U}^{\prime}$ of $X$ along $\mathrm{Im}(\sigma_{i})$ has exponent $e$ in the sense of Definition 3.4 under some (and hence, any) identification $\mathscr{U}^{\prime}\cong\mathscr{U}$.
4. Frobenius pull-back of parabolic bundles
In this section, we introduce the Frobenius pull-back of a parabolic bundle, generalizing the usual definition for the non-parabolic case. The main theorem of this section shows that this operation yields a bijective correspondence between parabolic bundles on $\mathscr{X}^{(N)}$ and parabolic $p^{N}$-flat bundles on $\mathscr{X}$ (cf. Theorem 4.5).
4.1. Parabolic Frobenius pull-back in a local situation I
As a first step, we consider the case where the underlying space is the $N$-th Frobenius twist $\mathscr{U}^{(N)}:=(U^{(N)},\sigma_{\circ}^{(N)})$ of $\mathscr{U}=(U,\sigma_{\circ})$ (cf. (3.14)) and $S$ is connected.
Let $m_{\circ}$ be a positive integer and $a_{\circ}:=(a_{\circ}^{[1]},\cdots,a_{\circ}^{[m_{\circ}]})$ an element of $\Xi_{m_{\circ},N}^{\leq}$. We shall set
$a_{\circ}/p^{N}:=(a_{\circ}^{[1]}/p^{N},\cdots,a_{\circ}^{[m_{\circ}]}/p^{N}).$ (4.1)
Just as in the case of $\mathscr{X}$, both the categories $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$ and $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}$ for this local situation can be defined in the same manner.
First, we construct an object of $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}$ by using an object of $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$. Let us take a parabolic bundle $\mathscr{E}:=(\mathcal{E},\mathfrak{e}_{\circ},a_{\circ}/p^{N})$ classified by $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$; the quasi-parabolic structure $\mathfrak{e}_{\circ}$ is given by
$\mathfrak{e}_{\circ}:\sigma_{\circ}^{*}\mathcal{E}=\mathcal{E}_{\circ}^{[m_{\circ}]}\xrightarrow{\mathfrak{e}_{\circ}^{[m_{\circ}]}}\mathcal{E}_{\circ}^{[m_{\circ}-1]}\xrightarrow{\mathfrak{e}_{\circ}^{[m_{\circ}-1]}}\cdots\xrightarrow{\mathfrak{e}_{\circ}^{[2]}}\mathcal{E}_{\circ}^{[1]}\xrightarrow{\mathfrak{e}_{\circ}^{[1]}}\mathcal{E}_{\circ}^{[0]}=0.$ (4.2)
For simplicity, we write $\mathcal{G}^{[j]}:=\sigma_{\circ*}\mathcal{E}_{\circ}^{[j]}$ ($j=0,\cdots,m_{\circ}$). For $j\in 1,\cdots,m_{\circ}$, the natural surjection $\varpi_{\mathscr{E}}^{[j]}:\mathcal{E}\twoheadrightarrow\mathcal{G}^{[j]}$ induces, via pull-back along $F^{(N)}_{U/S}$, a surjective morphism of $\mathcal{D}^{(N-1)}$-modules $F^{(N)*}_{U/S}\varpi_{\mathscr{E}}^{[j]}:(F^{(N)*}_{U/S}\mathcal{E},\nabla_{\mathcal{E}}^{\mathrm{can}})\twoheadrightarrow(F^{(N)*}_{U/S}\mathcal{G}^{[j]},\nabla_{\mathcal{G}^{[j]}}^{\mathrm{can}})$. Then, we obtain a composite surjection between $\mathcal{D}^{(N-1)}$-modules
$\widetilde{\varpi}_{\mathscr{E}}^{[j]}:$ $\ ((F^{(N)*}_{U/S}\mathcal{E})(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{E},p^{N}\cdot D_{\circ}}^{\mathrm{can}})$ (4.3)
$\twoheadrightarrow((F^{(N)*}_{U/S}\mathcal{G}^{[j]})(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{G}^{[j]},p^{N}\cdot D_{\circ}}^{\mathrm{can}})$
$\twoheadrightarrow((F^{(N)*}_{U/S}\mathcal{G}^{[j]})(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{G}^{[j]},p^{N}\cdot D_{\circ}}^{\mathrm{can}})\otimes(\mathcal{O}_{U}/\mathcal{O}_{U}((a^{[j]}_{\circ}-p^{N})\sigma_{\circ}),\overline{\nabla}_{0}),$
[…]
Its restriction to $U_{\circledast}:=U\setminus\mathrm{Im}(\sigma_{\circ})$ may be identified with the usual Frobenius pull-back of $\mathcal{E}$ with the canonical connection. This implies that $\nabla^{F}$ has vanishing $p^{N}$-curvature.
Moreover, we shall set
$\mathcal{E}^{F,[j]}_{\circ}:=\sigma^{*}_{\circ}\left(\mathcal{E}^{F}/(\mathcal{E}^{F}\cap\mathrm{Ker}(F^{(N)*}_{U/S}\varpi_{\mathscr{E}}^{[j]})(-D_{\circ}))\right).$ (4.5)
Then, there exists a sequence of natural surjections
$\mathfrak{e}^{F}_{\circ}:\mathcal{E}^{F,[m_{\circ}]}_{\circ}\xrightarrow{\mathfrak{e}_{\circ}^{F,[m_{\circ}]}}\mathcal{E}^{F,[m_{\circ}-1]}_{\circ}\xrightarrow{\mathfrak{e}_{\circ}^{F,[m_{i}-1]}}\cdots\xrightarrow{\mathfrak{e}_{\circ}^{F,[2]}}\mathcal{E}^{F,[1]}_{\circ}\xrightarrow{\mathfrak{e}_{\circ}^{F,[1]}}\mathcal{E}^{F,[0]}_{\circ}=0.$ (4.6)
Proposition 4.1.
The resulting collection
$\mathscr{E}_{\flat}^{F}:=(\mathcal{E}^{F},\nabla^{F},\mathfrak{e}^{F}_{\circ},a_{\circ})$ (4.7)
forms a parabolic $p^{N}$-flat bundle on $\mathscr{U}$ classified by $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$, and the equality $\mathcal{S}ol(\nabla^{F})=\mathcal{E}$ holds. Moreover, if $\ell_{\circ}:=(\ell_{\circ}^{[1]},\cdots,\ell_{\circ}^{[m_{\circ}]})$ denotes the type of $\mathfrak{e}_{\circ}$, then
$(\mathrm{det}(\mathcal{E}^{F}),\mathrm{det}(\nabla^{F}))\cong((F^{(N)*}_{U/S}\mathrm{det}(\mathcal{E}))(sD_{\circ}),\nabla^{\mathrm{can}}_{\mathrm{det}(\mathcal{E}),sD_{\circ}}),$ (4.8)
where $s:=\sum\limits_{j=1}^{m_{\circ}}a_{\circ}^{[j]}\ell^{[j]}_{\circ}$.
Proof.
We shall prove the first assertion. Let us set
$\mathcal{H}:=\bigoplus_{j^{\prime}=1}^{m_{\circ}}\mathcal{O}_{U^{(N)}}^{\oplus\ell_{\circ}^{[j^{\prime}]}}\hskip 14.22636pt\text{and}\hskip 14.22636pt\mathcal{H}_{\circ}^{[j]}:=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}\ (j=0,\cdots,m_{\circ}).$ (4.9)
In particular, we have $\sigma_{\circ}^{*}\mathcal{H}=\bigoplus_{j^{\prime}=1}^{m_{\circ}}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$. For each $j=1,\cdots,m_{\circ}$, denote by $\mathfrak{h}_{\circ}^{[j]}$ the projection $\mathcal{H}_{\circ}^{[j]}\twoheadrightarrow\mathcal{H}_{\circ}^{[j-1]}$ given by forgetting the last direct summand. The triple
$\mathscr{H}:=(\mathcal{H},\mathfrak{h}_{\circ},a_{\circ}/p^{N}),$ (4.10)
where $\mathfrak{h}_{\circ}:=\{\mathfrak{h}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$, specifies a parabolic bundle on $\mathscr{U}^{(N)}$. Since the assertion is of local nature, we can suppose that there exists an isomorphism $\tau:\mathscr{E}\xrightarrow{\sim}\mathscr{H}$. It follows that the problem can be reduced to the case of $\mathscr{E}=\mathscr{H}$.
For each $j=1,\cdots,m_{\circ}$, we have
$(\mathrm{Ker}(\widetilde{\varpi}_{\mathscr{H}}^{[j]}),\nabla_{\mathrm{Ker}}^{[j]})=\bigoplus_{j^{\prime}=1}^{j}(\mathcal{O}_{U}(a^{[j^{\prime}]}_{\circ}D_{\circ}),\nabla_{0})^{\oplus\ell_{\circ}^{[j^{\prime}]}}\oplus\bigoplus_{j^{\prime}=j+1}^{m_{\circ}}(\mathcal{O}_{U}(p^{N}D_{\circ}),\nabla_{0})^{\oplus\ell^{[j^{\prime}]}_{\circ}},$ (4.11)
where we abuse notation by writing $\nabla_{0}$ for the restrictions of the trivial $\mathcal{D}^{(N-1)}$-module structure on $\mathcal{O}_{U}$ to various $\mathcal{O}_{X}$-submodules. This implies
$\mathscr{H}^{F}=\bigoplus_{j^{\prime}=1}^{m_{i}}(\mathcal{O}_{U}(a^{[j^{\prime}]}_{\circ}D_{\circ}),\nabla_{0})^{\oplus\ell_{j^{\prime}}},\hskip 14.22636pt\mathcal{H}_{\circ}^{F[j]}=\bigoplus_{j^{\prime}=1}^{j}\sigma_{\circ}^{*}\left(\mathcal{O}_{U}(a^{[j^{\prime}]}_{\circ}D_{\circ})^{\oplus\ell_{j^{\prime}}}\right)$ (4.12)
($j=0,\cdots,m_{\circ}$), and the projection $\mathfrak{h}_{\circ}^{F[j]}$ is given by forgetting the last direct summand. Thus, $(\mathcal{E}^{F},\mathfrak{h}_{\circ}^{F})$ forms a quasi-parabolic bundle on $\mathscr{U}$. Moreover, since $(\mathcal{O}_{U}(a_{\circ}^{[j]}D_{\oslash}),\nabla_{0})\cong(\mathcal{O}_{U},\nabla_{a_{\circ}^{[j]}})$, the description of $\nabla_{\mathrm{Ker}}^{[j]}$ obtained in (4.11) and the equality (3.18) together imply that $\mathscr{H}^{F}_{\flat}:=(\mathcal{E}^{F},\nabla^{F},\mathfrak{h}_{\circ}^{F},a_{\circ})$ belongs to $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}$. This completes the proof of the first assertion.
The second assertion follows from the definition of $(\mathcal{E}^{F},\nabla^{F})$ together with the discussion in the proof of the first assertion. ∎
Note that the formation of $\mathscr{E}\mapsto\mathscr{E}^{F}_{\flat}$ is functorial. This is to say, if $h:\mathscr{E}\rightarrow\mathscr{E}^{\prime}$ is a morphism in $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$, then this yields naturally a morphism $h^{F}:\mathscr{E}^{F}_{\flat}\rightarrow{\mathscr{E}^{\prime}_{\flat}}^{F}$. Thus, the assignments $\mathscr{E}\mapsto\mathscr{E}_{\flat}^{F}$ and $h\mapsto h^{F}$ together define a functor $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}\rightarrow\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$.
4.2. Parabolic Frobenius pull-back in a local situation II
Next, we give an inverse direction of the above construction. Let $\mathscr{F}_{\flat}:=(\mathcal{F},\nabla,\mathfrak{f}_{\circ},a_{0})$ (where $\mathscr{F}:=(\mathcal{F},\mathfrak{f}_{\circ})$, $\mathfrak{f}_{\circ}:=\{\mathfrak{f}_{\circ}^{[j]}\}_{j=1}^{m_{\circ}}$) be a parabolic $p^{N}$-flat bundle on $\mathscr{U}$ classified by $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$. We shall set $\mathcal{F}^{\nabla}:=\mathcal{S}ol(\nabla)$, which specifies a rank-$n$ vector bundle on $U^{(N)}$. For each $j=0,\cdots,m_{\circ}$, $\nabla$ restricts to a $\mathcal{D}^{(N-1)}$-module structure $\nabla^{[j]}$ on $\mathrm{Ker}(\varpi_{\mathscr{F}}^{[j]})(-(a_{\circ}^{[j]}+1)D_{\circ})$. Then, $\mathcal{S}ol(\nabla^{[j]})$ specifies a vector bundle on $U^{(N)}$ equipped with a natural inclusion $\mathcal{S}ol(\nabla^{[j]})\hookrightarrow\mathcal{S}ol(\nabla)$, which becomes an isomorphism when restricted to $U^{(N)}\setminus\mathrm{Im}(\sigma_{\circ}^{(N)})$. The $\mathcal{O}_{S}$-modules $\mathcal{F}^{\nabla[j]}_{\circ}:=\sigma_{\circ}^{(N)*}(\mathcal{S}ol(\nabla)/\mathcal{S}ol(\nabla^{[j]}))$ fit into a sequence of surjections
$\mathfrak{f}_{\circ}^{\nabla}:\sigma_{\circ}^{(N)*}\mathcal{F}^{\nabla}=\mathcal{F}_{\circ}^{\nabla[m_{\circ}]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[m_{\circ}]}}\mathcal{F}_{\circ}^{\nabla[m_{\circ}-1]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[m_{\circ}-1]}}\cdots\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[2]}}\mathcal{F}_{\circ}^{\nabla[1]}\xrightarrow{\mathfrak{f}_{\circ}^{\nabla[1]}}\mathcal{F}_{\circ}^{\nabla[0]}=0.$ (4.13)
Then, we can prove the following assertion.
Proposition 4.2.
  • (i)
    The collection
    $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\mathfrak{f}^{\nabla}_{\circ},a_{\circ}/p^{N})$ (4.14)
    forms a parabolic bundle on $\mathscr{U}^{(N)}$ classified by $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$.
  • (ii)
    Let $\mathscr{E}$ and $\mathscr{F}_{\flat}$ be objects in $\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$ and $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$, respectively. Then, there exist canonical isomorphisms
    $\mathscr{E}\xrightarrow{\sim}(\mathscr{E}_{\flat}^{F})^{\nabla},\hskip 14.22636pt(\mathscr{F}^{\nabla})^{F}_{\flat}\xrightarrow$$\sim$$\mathscr{F}_{\flat}$ (4.15)
[…]
Proof.
First, we shall prove assertion (i). Denote by $\ell_{\circ}:=(\ell_{\circ}^{[1]},\cdots,\ell_{\circ}^{[m_{\circ}]})$ the type of $\mathfrak{f}_{\circ}$. The assertion is of local nature, so $\mathscr{F}:=(\mathcal{F},\nabla)$ may be identified with a direct sum of $\mathscr{O}_{d,\flat}$’s (cf. the discussion around (3.19)). Since $(\mathcal{F},\nabla)$ has parabolic weights $a_{0}$, the equality (3.18) implies that the problem is reduced to the situation where $\mathscr{F}$ satisfies the following conditions:
  • $\mathscr{F}$ coincides with the direct sum
    $\mathscr{F}=\bigoplus_{j=1}^{m_{\circ}}(\mathcal{O}_{U}(a_{\circ}^{[j]}D_{\circ}),\nabla_{0})^{\oplus\ell_{\circ}^{[j]}}$ (4.16)
    where, just as in the proof of Proposition 4.1, we abuse notation by writing $\nabla_{0}$ for various restrictions of the trivial one;
  • For each $j=0,\cdots,m_{\circ}$, the equality $\mathcal{F}_{\circ}^{[j]}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds. Also, the projection $\mathfrak{f}_{\circ}^{[j]}$ coincides with the projection $\mathcal{F}_{\circ}^{[j]}$ given by forgetting the last direct summand.
Then, it is verified that $\mathcal{F}^{\nabla}=\mathcal{S}ol(\nabla)=\bigoplus_{j=1}^{m_{\circ}}\mathcal{O}_{U^{(N)}}^{\oplus\ell_{\circ}^{[j]}}$ and
$\mathcal{S}ol(\nabla^{[j]})=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{U^{(N)}}(-D_{\circ}^{(N)})^{\oplus\ell_{\circ}^{[j^{\prime}]}}\oplus\bigoplus_{j^{\prime}=j+1}^{m_{\circ}}\mathcal{O}_{U^{(N)}}$ (4.17)
($j=0,\cdots,m_{\circ}$), where $D^{(N)}_{\circ}:=\mathrm{Im}(\sigma_{\circ}^{(N)})$. Hence, the equality $\mathcal{F}^{\nabla[j]}_{\circ}=\bigoplus_{j^{\prime}=1}^{j}\mathcal{O}_{S}^{\oplus\ell_{\circ}^{[j^{\prime}]}}$ holds and $\mathfrak{f}_{\circ}^{\nabla[j]}:\mathcal{F}^{\nabla[j]}_{\circ}\twoheadrightarrow\mathcal{F}^{\nabla[j-1]}_{\circ}$ coincides the projection given by forgetting the last direct summand. Thus, $\mathscr{F}^{\nabla}$ specifies an object in $\mathcal{B}un_{\mathscr{U}^{(N)},a_{0}/p^{N}}$, and this completes the proof of assertion (i).
Next, we shall prove assertion (ii). By considering the various definitions involved, one can construct the morphisms displayed in (4.15). To prove that these morphisms are isomorphisms, we may assume that $S$ is the spectrum of a local ring. Thus, it suffices to consider the case where $\mathscr{E}$ and $\mathscr{F}_{\flat}$ are of the forms (4.10) and (4.16), respectively, and the assertion can be proved immediately from the discussions in the proofs of Propositions 4.1 and 4.2, (i). ∎
Since the formation of $\mathscr{F}_{\flat}\mapsto\mathscr{F}^{\nabla}$ is verified to be functorial, this operation yields a functor $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}\rightarrow\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}$.
Theorem 4.3.
The assignments $\mathscr{E}\mapsto\mathscr{E}^{F}_{\flat}$ and $\mathscr{F}_{\flat}\mapsto\mathscr{F}^{\nabla}$ together determine an equivalence of categories
$\mathcal{B}un_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}.$ (4.18)
Proof.
The assertion follows from Proposition 4.2, (ii). ∎
4.3. Parabolic Frobenius pull-back in a global situation
As the final step of formulating the Frobenius pull-back, let us consider the case where underlying space is the globally defined curve $\mathscr{X}$. We may still assume that $S$ is connected.
Let us choose an $r$-tuple of positive integers $(m_{1},\cdots,m_{r})$ and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$, where $a_{i}:=(a_{i}^{[1]},\cdots,a_{i}^{[m_{i}]})$ ($i=1,\cdots,r$). We shall set
$a_{i}/p^{N}:=(a_{i}^{[1]}/p^{N},\cdots,a_{i}^{[m_{i}]}/p^{N}),\hskip 22.76219pt\vec{a}/p^{N}:=\left(a_{1}/p^{N},\cdots,a_{r}/p^{N}\right).$ (4.19)
Also, let $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ be a parabolic bundle on $\mathscr{X}^{(N)}$ belonging to $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$. We apply the operation $(-)_{\flat}^{F}$ obtained in § 4.1 to the restriction of $\mathscr{E}$ to the formal neighborhood of $\sigma_{i}$ for every $i=1,\cdots,r$. By gluing together the resulting parabolic Frobenius pull-backs and the usual Frobenius pull-back over $X\setminus\bigcup_{i=1}^{r}\mathrm{Im}(\sigma_{i})$, we obtain a parabolic $p^{N}$-flat bundle
$\mathscr{E}_{\flat}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ (4.20)
on $\mathscr{X}$, which belongs to $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}$. If $(\ell_{i}^{[1]},\cdots,\ell_{i}^{[m_{i}]})$ denotes the type of $\vec{\mathfrak{e}}$ at $\sigma_{i}$, then the second assertion in Proposition 4.1 implies that $\mathcal{S}ol(\nabla^{F})=\mathcal{E}$ and
$(\mathrm{det}(\mathcal{E}^{F}),\mathrm{det}(\nabla^{F}))\cong((F^{(N)*}_{X/S}\mathrm{det}(\mathcal{E}))(D_{\star}),\nabla^{\mathrm{can}}_{\mathrm{det}(\mathcal{E}),D_{\star}}),$ (4.21)
where $D_{\star}:=\sum\limits_{i=1}^{r}\left(\sum\limits_{j=1}^{n_{i}}a_{i}^{[j]}\ell_{i}^{[j]}\right)\sigma_{i}$. This implies
$\mathrm{deg}(\mathcal{E}^{F})=p^{N}\cdot\mathrm{par}\text{-}\mathrm{deg}(\mathscr{E}),\hskip 14.22636pt\mu(\mathcal{E}^{F})=p^{N}\cdot\mathrm{par}\text{-}\mu(\mathscr{E}),$ (4.22)
where $\mathrm{deg}(-)$ and $\mu(-)$ denote the usual (i.e., non-parabolic) degree and slope, respectively.
On the other hand, let us take a parabolic $p^{N}$-flat bundle $\mathscr{F}_{\flat}$ classified by $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}$. Similarly to the above discussion, we can apply the operation $(-)^{\nabla}$ obtained in § 4.2 to obtain, from $\mathscr{F}_{\flat}$, a parabolic bundle
$\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathfrak{f}}^{\,\nabla},\vec{a}/p^{N})$ (4.23)
on $\mathscr{X}^{(N)}$ belonging to $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$.
math 4.4.
[…]
The following theorem can be proved immediately by combining the higher-level generalization of Cartier’s theorem by B. Le Stum and A. Quirós (cf.  [Kat, Theorem 5.1],  [LeQu, Corollary 3.2.4]) and Theorem 4.3 (or the second assertion of Proposition 4.2, (ii)) applied to the formal neighborhoods of the marked points.
Theorem 4.5 (cf. Theorem A).
Let $(m_{1},\cdots,m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an element of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$. Then, the assignments $\mathscr{E}\mapsto\mathscr{E}^{F}_{\flat}$ and $\mathscr{F}_{\flat}\mapsto\mathscr{F}^{\nabla}$ together give an equivalence of categories
$\mathcal{B}un_{\mathscr{X}^{(N-1)},\vec{a}/p^{N}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}.$ (4.24)
Moreover, the formation of this equivalence is functorial with respect to $S$.
In particular, we obtain the following assertion.
Theorem 4.6.
Let us keep the notation in the previous theorem. Also, let $\mathscr{E}$ be a parabolic bundle classified by $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$. Then, for a parabolic subbundle $\mathscr{G}$ of $\mathscr{E}$, its Frobenius pull-back $\mathscr{G}^{F}_{\flat}$ forms a parabolic $p^{N}$-flat subbundle of $\mathscr{E}^{F}_{\flat}$. Moreover, the resulting assignment $\mathscr{G}\mapsto\mathscr{G}^{F}_{\flat}$ gives a bijective correspondence between parabolic subbundles of $\mathscr{E}$ and parabolic $p^{N}$-flat subbundles of $\mathscr{E}^{F}_{\flat}$.
Proof.
The assertions follow from the various definitions involved. ∎
4.4. Transitivity of parabolic Frobenius pull-backs
[…]
For each integer $a$ in $\{0,\cdots,p^{N}-1\}$, let $(s_{1}^{M}(a),s_{2}^{M}(a))$ be the unique pair of integers satisfying that $a=s_{1}^{M}(a)+p^{M}\cdot s_{2}^{M}(a)$ and $0\leq s_{1}^{M}(a)<p^{M}$, $0\leq s_{2}^{M}(a)<p^{N-M}$. We choose an $r$-tuple of positive integers $(m_{1},\cdots,m_{r})$ and an element $\vec{a}:=(a_{1},\cdots,a_{r})$, where $a_{i}=(a_{i}^{[1]},\cdots,a_{i}^{[m_{i}]})$, of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ satisfying
$s_{1}^{M}(a_{i}^{[1]})\leq\cdots\leq s_{1}^{M}(a_{i}^{[m_{i}]})\hskip 14.22636pt\text{and}\hskip 14.22636pts_{2}^{M}(a_{i}^{[1]})\leq\cdots\leq s_{2}^{M}(a_{i}^{[m_{i}]})$ (4.25)
for every $i=1,\cdots,r$. By putting $s_{t}^{M}(a_{i}):=(s_{t}^{M}(a_{i}^{[1]}),\cdots,s_{t}^{M}(a_{i}^{[m_{i}]}))$ ($t\in\{1,2\}$, $i\in\{1,\cdots,r\}$), we obtain a collection of parabolic weights $s_{t}^{M}(\vec{a}):=(s_{t}^{M}(a_{1}),\cdots,s_{t}^{M}(a_{r}))$.
Now, let $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ be a parabolic bundle classified by $\mathcal{B}un_{\mathscr{X}^{(N)},\vec{a}/p^{N}}$; it associates the parabolic Frobenius pull-back $\mathcal{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ by $F_{X/S}^{(N)}$. Also, since $s_{2}^{M}(\vec{a})$ belongs to $\prod_{i=1}^{r}\Xi_{m_{i},N-M}^{\leq}$, one obtains the parabolic Frobenius pull-back of $(\mathcal{E},\vec{\mathfrak{f}},s_{2}^{M}(\vec{a})/p^{N-M})$ by $F_{X^{(M)}/S}^{(N-M)}:X^{(M)}\rightarrow X^{(N)}$, which we denote by $\mathcal{E}^{\prime}:=(\mathcal{E}^{\prime},\nabla^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{2}^{M}(\vec{a}))$. The resulting triple $(\mathcal{E}^{\prime},\vec{\mathfrak{e}}\,^{\prime},s_{1}^{M}(\vec{a})/p^{M})$ specifies an object in the category $\mathcal{B}un_{\mathscr{X}^{(M)},s_{1}^{M}(\vec{a})/p^{M}}$, so it associates the parabolic Frobenius pull-back by $F_{X/S}^{(M)}$; we denote it by $\mathscr{E}^{\prime\prime}:=(\mathcal{E}^{\prime\prime},\nabla^{\prime\prime},\vec{\mathfrak{e}}\,^{\prime\prime},s_{1}^{M}(\vec{a}))$.
It follows from  [Mon, Corollaire 3.3.1] that the Frobenius pull-back of $\nabla^{\prime}$ (in the usual sense) gives a $\mathcal{D}^{(N-1)}$-module structure on $F^{(M)*}_{X/S}\mathcal{E}^{\prime}$, and the definition of parabolic Frobenius pull-back implies that this extends to a $\mathcal{D}^{(N-1)}$-module structure $\nabla^{F^{\prime\prime}}$ on $\mathcal{E}^{\prime\prime}$. By construction, the $\mathcal{D}^{(M-1)}$-module structure on $\mathcal{E}^{\prime\prime}$ obtained from $\nabla^{F^{\prime\prime}}$ by reducing the level coincides with $\nabla^{\prime\prime}$. Moreover, the following proposition can be proved by the various definitions involved.
Proposition 4.7.
The natural isomorphism $F^{(M)*}_{X/S}F_{X/S}^{(N-M)*}\mathcal{E}\xrightarrow{\sim}F^{(N)*}_{X/S}\mathcal{E}$ extends to an isomorphism of parabolic $p^{N}$-flat bundles
$(\mathcal{E}^{\prime\prime},\nabla^{F^{\prime\prime}},\vec{\mathfrak{e}}\,^{\prime\prime},\vec{a})\xrightarrow{\sim}(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}},\vec{a}).$ (4.26)
4.5. Parabolic Higgs bundles
The celebrated Ogus-Vologodsky correspondence (cf.  [OgVo]) generalizes Cartier’s theorem in such a way that involves nilpotent Higgs bundles, and there exist some variants including higher-level $\mathcal{D}$-modules (cf.  [GLQ],  [Ohk],  [Sch]). In the rest of this section, we will give another slight extension, i.e., the Ogus-Vologodsky correspondence for parabolic Higgs bundles, extending the equivalence asserted in Theorem 4.5.
Suppose here that we are given a lifting $\widetilde{S}$ of $S$ modulo $p^{2}$ and a lifting $\widetilde{\mathscr{X}}^{(N)}:=(\widetilde{X}^{(N)},\{\widetilde{\sigma}_{i}\}_{i=1}^{r})$ of $\mathscr{X}^{(N)}$ over $\widetilde{S}$. We write $\mathscr{X}_{\emptyset}$ for the unpointed smooth curve $(X,\emptyset)$. Recall from  [Ohk] that the lifting $\widetilde{X}/\widetilde{S}$ associates a canonical equivalence of categories between the following categories:
  • (a)
    the category of $\mathcal{O}_{X^{(N)}}$-modules $\mathcal{E}$ equipped with a Higgs field $\theta:\mathcal{E}\rightarrow\Omega_{X^{(N)}/S}\otimes\mathcal{E}$ with $\theta^{p-1}=0$;
  • (b)
    the category of $\mathcal{D}_{X/S}^{(N-1)}$-modules $(\mathcal{F},\nabla)$ whose $p^{N}$-curvature $\psi_{(\mathcal{F},\nabla)}$ satisfies $\psi_{(\mathcal{F},\nabla)}^{p-1}=0$.
Note that this equivalence can be formulated in the case where $\mathscr{X}$ is replaced with the local curve $\mathscr{U}$ defined as in (3.15) (after fixing a lifting of $\mathscr{U}^{(N)}$ modulo $p^{2}$). For each object $(\mathcal{E},\theta)$ (resp., $(\mathcal{F},\nabla)$) in the category (a) (resp., (b)), the counterpart via the above equivalence will be denoted by $C^{-1}_{\widetilde{\mathscr{X}}^{(N)}}(\mathcal{E},\theta)$ (resp., $C_{\widetilde{\mathscr{X}}^{(N)}}(\mathcal{F},\nabla)$).
Definition 4.8.
  • (i)
    A non-logarithmic parabolic Higgs bundle on $\mathscr{X}$ is a collection of data
    $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha},\theta)$ (4.27)
[…]
    Also, one can define the notion of a morphism between two non-logarithmic parabolic Higgs bundles.
  • (ii)
    We say that a non-logarithmic parabolic Higgs bundle $\mathscr{F}_{\sharp}:=(\mathcal{F},\vec{\mathfrak{f}},\vec{\alpha},\theta)$ is $p^{N}$-nilpotent if $\theta^{p-1}=0$.
Let $(m_{1},\cdots m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an element of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$. Denote by
$\mathcal{H}ig_{\mathscr{X},\vec{a}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$ (4.28)
the category of $p^{N}$-nilpotent non-logarithmic parabolic Higgs bundles on $\mathscr{X}$ of parabolic weights $\vec{a}/p^{N}$. Since each vector bundle may be identified with a Higgs bundle with vanishing Higgs field, $\mathcal{H}ig_{\mathscr{X},\vec{a}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$ contains $\mathcal{B}un_{\mathscr{X},\vec{a}/p^{N}}$ as a full subcategory. The notion introduced in Definition 4.8 can be formulated when the underlying curve is replaced with $\mathscr{U}$. In particular, we have the category $\mathcal{H}ig_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$ for parabolic weights $a_{\circ}/p^{N}$.
[…]
$\widetilde{\varpi}^{[j]}_{\mathscr{E}}$ $:(\mathcal{H}(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}})$ (4.29)
$\ \ \twoheadrightarrow(\mathcal{H}_{i}^{[j]}(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}}^{[j]})$
$\ \ \twoheadrightarrow(\mathcal{H}_{i}^{[j]}(p^{N}\cdot D_{\circ}),\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}}^{[j]})\otimes(\mathcal{O}_{X}/\mathcal{O}_{X}((a_{\circ}^{[j]}-p^{N})\overline{\sigma}_{i}),\overline{\nabla}_{0})$ (4.30)
where
  • $\nabla_{\mathcal{H},p^{N}\cdot D_{\circ}}$ (resp., $\nabla^{[j]}_{\mathcal{H},p^{N}\cdot D_{\circ}}$) denotes the $\mathcal{D}^{(N-1)}_{\circ}$-module structure on $\mathcal{H}(p^{N}\cdot D_{\circ})$ (resp., $\mathcal{H}^{[j]}(p^{N}\cdot D_{\circ})$) extending $\nabla_{\mathcal{H}}$ (resp., $\nabla^{[j]}_{\mathcal{H}}$);
  • the first arrow arrow is obtained by extending $\check{\varpi}_{\mathscr{E}}^{[j]}$, and the second arrow arises from the natural quotient $(\mathcal{O}_{U},\nabla_{0})\twoheadrightarrow(\mathcal{O}_{U}/\mathcal{O}_{U}((a_{\circ}^{[j]}-p^{N})\overline{\sigma}_{\circ}),\overline{\nabla}_{0})$.
The kernel of this composite determines a $\mathcal{D}^{(N-1)}_{\circ}$-module $(\mathrm{Ker}(\widetilde{\varpi}_{\mathscr{E}}^{[j]}),\nabla_{\mathrm{Ker}}^{[j]})$. Applying this construction to various $j$’s and then taking their intersection, we obtain a $\mathcal{D}_{\circ}^{(N-1)}$-module
$\mathscr{E}^{F}=(\mathcal{E}^{F},\nabla^{F}):=\bigcap_{j=1}^{m_{\circ}}(\mathrm{Ker}(\widetilde{\varpi}_{\mathscr{E}}^{[j]}),\nabla_{\mathrm{Ker}}^{[j]}).$ (4.31)
Moreover, just as in (4.5) and (4.6), we obtain a quasi-parabolic structure $\mathfrak{e}^{F}_{\circ}$ on $\mathcal{E}^{F}$. The resulting collection
$\mathscr{E}^{F}_{\flat}:=(\mathcal{E}^{F},\nabla^{F},\mathfrak{e}_{\circ},a_{\circ})$ (4.32)
forms a parabolic $p^{N}$-nilpotent bundle on $\mathscr{U}$. (This construction for vanishing Higgs fields coincides with that in (4.20).)
Definition 4.9.
  • (i)
    Let $\mathscr{F}_{\flat}$ be a parabolic $p^{N}$-nilpotent bundle on $\mathscr{U}$ of parabolic weights $a_{\circ}$. We say that $\mathscr{F}_{\flat}$ is of restricted type if it is isomorphic to $\mathscr{E}^{F}_{\flat}$ for some object $\mathscr{E}$ in $\mathcal{H}ig_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}$.
  • (ii)
    Let $\mathscr{F}_{\flat}$ be a parabolic $p^{N}$-nilpotent bundle on $\mathscr{X}$. We say that $\mathscr{F}_{\flat}$ is of restricted type if, for each $i=1,\cdots,r$, its restriction to the formal completion $\mathscr{U}^{\prime}$ of $X$ along $\mathrm{Im}(\sigma_{i})$ is, locally on $S$, of restricted type under some (and hence, any) identification $\mathscr{U}^{\prime}\cong\mathscr{U}$.
Denote by
$\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\mathrm{res}}\ \left(\text{resp.,}\ \mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\mathrm{res}}\right)$ (4.33)
the full subcategory of $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\mathrm{nilp}}$ (resp., $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\mathrm{nilp}}$) consisting of parabolic $p^{N}$-nilpotent bundles of restricted type; it contains $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\psi=0}$ (resp., $\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\psi=0}$) as a full subcategory.
4.6. Parabolic generalization of O-V correspondence
Let $\mathscr{E}_{\sharp}$ and $\mathscr{E}^{F}_{\flat}$ be as in the previous subsection. Denote by $\mathcal{F}$ (resp., $\mathcal{F}^{[j]}$) the maximal element (with respect to the inclusion relation) among all the $\mathcal{D}_{\circ}^{(N-1)}$-submodules $\mathcal{F}^{\prime}$ of $\mathcal{E}^{F}$ (resp., $\mathrm{Ker}(\varpi^{[j]}_{\mathscr{E}^{F}})(-(a_{\circ}^{[j]}+1)D_{\circ})$) whose $\mathcal{D}_{\circ}^{(N-1)}$-actions extend to $\mathcal{D}_{U/S}^{(N-1)}$-actions. By definition, $\mathcal{F}$ (resp., $\mathcal{F}^{[j]}$) is equipped with a $\mathcal{D}_{U/S}^{(N-1)}$-module structure $\nabla_{\mathcal{F}}$ (resp., $\nabla_{\mathcal{F}^{[j]}}$). By applying the functor $C_{\widetilde{\mathscr{U}}^{(N)}}(-)$ to $(\mathcal{F},\nabla_{\mathcal{F}})$ (resp., $(\mathcal{F}^{[j]},\nabla_{\mathcal{F}^{[j]}})$), we obtain a Higgs bundle $(\mathcal{E}_{\star},\theta_{\star})$ (resp., $(\mathcal{E}_{\star}^{[j]},\theta_{\star}^{[j]})$). The $\mathcal{O}_{S}$-module $\mathcal{E}_{\star,\circ}^{[j]}:=\sigma_{\circ}^{(N)*}(\mathcal{E}_{\star}/\mathcal{E}_{\star}^{[j]})$ fit into a sequence of surjections
$\mathfrak{e}_{\star,\circ}:\sigma_{\circ}^{(N)*}\mathcal{E}_{\star}=\mathcal{E}_{\star,\circ}^{[m_{\circ}]}\xrightarrow{\mathfrak{e}_{\star,\circ}^{[m_{\circ}]}}\mathcal{E}_{\star,\circ}^{[m_{\circ}-1]}\xrightarrow{\mathfrak{e}_{\star,\circ}^{[m_{\circ}-1]}}\cdots\xrightarrow{\mathfrak{e}_{\star,\circ}^{[2]}}\mathcal{E}_{\star,\circ}^{[1]}\xrightarrow{\mathfrak{e}_{\star,\circ}^{[0]}}\mathcal{E}_{\star,\circ}^{[0]}=0.$ (4.34)
Thus, we have obtained a collection
$\mathscr{E}_{\star,\sharp}:=(\mathcal{E}_{\star},\theta_{\star},\mathfrak{e}_{\star,\circ},a_{\circ}/p^{N}).$ (4.35)
Lemma 4.10.
$\mathscr{E}_{\star,\sharp}$ forms a non-logarithmic $p^{N}$-nilpotent Higgs bundle, and there exists an isomorphism $\mathscr{E}_{\sharp}\xrightarrow{\sim}\mathscr{E}$$_{\star,\sharp}$ functorial with respect to $\mathscr{E}_{\sharp}$. In particular, $\mathscr{E}_{\sharp}\mapsto\mathscr{E}^{F}_{\flat}$ defines an equivalence of categories
$\mathcal{H}ig_{\mathscr{U}^{(N)},a_{\circ}/p^{N}}^{\mathrm{non}\text{-}\mathrm{log}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{U},a_{\circ}}^{\mathrm{res}}.$ (4.36)
Proof.
[…]
Now, let us go back to the global situation. One can glue together the equivalence between the categories (a) and (b) (mentioned above) with $X$ replaced by $X\setminus\bigcup_{i=1}^{r}\mathrm{Im}(\sigma_{i})$ and the equivalence (4.36) in the case where $\mathscr{U}$ is taken to be the formal completion along every marked point. As a result, we obtain the following assertion.
Theorem 4.11.
There exists a canonical equivalence of categories
$\mathcal{H}ig_{\mathscr{X}^{(N)},\vec{a}/p^{N}}^{\mathrm{non\text{-}log}}\xrightarrow{\sim}\mathcal{D}^{(N-1)}\text{-}\mathcal{B}un_{\mathscr{X},\vec{a}}^{\mathrm{res}}$ (4.37)
such that the following square diagram is $1$-commutative:
Moreover, the formation of this equivalence is functorial with respect to $S$.
5. Dormant $\mathrm{GL}_{n}^{(N)}$-opers
In this section, we recall from  [Wak4] the definition of a dormant $\mathrm{GL}_{n}^{(N)}$-oper. Roughly speaking, a dormant $\mathrm{GL}_{n}^{(N)}$-oper is a $p^{N}$-flat bundle equipped with a complete flag satisfying a strict form of Griffiths transversality. We will prove in the next section that these objects defined on $\mathscr{X}$ correspond bijectively to certain parabolic bundles on $\mathscr{X}^{(N)}$ destabilized by parabolic Frobenius pull-back.
In the rest of the present paper, we assume that the pair of nonnegative integers $(g,r)$ satisfies 2g2+r>02g-2+r>0.
5.1. Definition of a dormant $\mathrm{GL}_{n}^{(N)}$-oper
Let $n$ be a positive integer. Consider a collection of data
$\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n}),$ (5.1)
where
  • $\mathcal{F}$ is a vector bundle on $X$ of rank $n$;
  • $\nabla$ is a $\mathcal{D}^{(N-1)}$-module structure on $\mathcal{F}$;
  • $\{\mathcal{F}^{j}\}_{j=0}^{n}$ is an $n$-step decreasing filtration
    $0=\mathcal{F}^{n}\subseteq\mathcal{F}^{n-1}\subseteq\dotsm\subseteq\mathcal{F}^{0}=\mathcal{F}$ (5.2)
    on $\mathcal{F}$ consisting of subbundles such that the subquotients $\mathcal{F}^{j}/\mathcal{F}^{j+1}$ are line bundles.
Definition 5.1.
  • (i)
    We shall say that $\mathscr{F}^{\heartsuit}$ is a $\mathrm{GL}_{n}^{(N)}$-oper (or a $\mathrm{GL}_{n}$-oper of level $N$) on $\mathscr{X}$ if, for every $j=0,\cdots,n-1$, the $\mathcal{O}_{X}$-linear morphism $\mathcal{D}^{(N-1)}\otimes\mathcal{F}\rightarrow\mathcal{F}$ induced by $\nabla$ restricts to an isomorphism
    $\mathcal{D}_{<n-j}^{(N-1)}\otimes\mathcal{F}^{n-1}\xrightarrow{\sim}\mathcal{F}^{j}.$ (5.3)
    The notion of an isomorphism between $\mathrm{GL}_{n}^{(N)}$-opers can be defined in a natural manner.
  • (ii)
    We shall say that a $\mathrm{GL}_{n}^{(N)}$-oper $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j})$ is dormant if $\nabla$ has vanishing $p^{N}$-curvature.
Let $\vec{a}:=(a_{1},\cdots,a_{r})$ (where $a_{i}:=(a_{i}^{[1]},\cdots,a_{i}^{[n]})$) be an element of $(\Xi_{n,N}^{\leq})^{\times r}$, and let $\mathcal{L}_{\flat}:=(\mathcal{L},\nabla_{\mathcal{L}})$ be a $p^{N}$-flat line bundle on $\mathscr{X}$, i.e., a pair consisting of a line bundle $\mathcal{L}$ on $X$ and a $\mathcal{D}^{(N-1)}$-module structure $\nabla_{\mathcal{L}}$ on $\mathcal{L}$ with vanishing $p^{N}$-curvature. For each $S$-scheme $s:S^{\prime}\rightarrow S$, we shall write
$\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}(S^{\prime})$ (5.4)
for the groupoid classifying collections $(\mathscr{F}^{\heartsuit},\eta)$, where
  • $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$ is a dormant $\mathrm{GL}^{(N)}_{n}-$oper on $\mathscr{X}_{S^{\prime}}$ whose exponent at $\sigma_{i}$ coincides with the multiset $a^{[1]}_{i},\cdots,a^{[n]}_{i}]$ for every $i=1,\cdots,r$;
  • $\eta$ denotes an isomorphism of $\mathcal{D}^{(N-1)}_{X^{\mathrm{log}}_{S^{\prime}}/S^{\prime}}-$modules $\mathrm{det}(\mathcal{F},\nabla)\xrightarrow{\sim}s^{*}\mathcal{L}_{\flat}$, where $s^{*}\mathcal{L}_{\flat}$ denotes the base-change of $\mathcal{L}_{\flat}$ along $s$.
Then, we obtain the category fibered in groupoids
$\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ (5.5)
over $\mathcal{S}ch_{/S}$ whose fiber over each $S$-scheme $S^{\prime}$ is $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}(S^{\prime})$. It is immediately verified that $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ forms a (possibly empty) stack over $S$. Since the determinants of $\mathrm{GL}_{n}^{(N)}$-opers in $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ are isomorphic to $\mathcal{L}_{\flat}$, this stack becomes empty unless the exponent of $\mathcal{L}_{\flat}$ at $\sigma_{i}$ coincides with the sum $\sum\limits_{j=1}^{n}a_{i}^{[j]}$ modulo $p^{N}$ (for every $i=1,\cdots,r$). Also, it follows from  [Wak4, Proposition 6.3.5] that this stack becomes empty unless $a_{i}$ lies in $\Xi_{n,N}^{<}\left(\subseteq\Xi_{n,N}^{\leq}\right)$ for every $i$.
Remark 5.2.
Suppose that we are given an object $(\mathscr{F}^{\heartsuit},\eta)$ of $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$, where $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$. Since (5.3) is an isomorphism, there exists a composite isomorphism
$\mathcal{F}^{j}/\mathcal{F}^{j+1}$ $\xrightarrow{\sim}(\mathcal{D}_{<n-j}^{(N-1)}\otimes\mathcal{F}^{n-1})/(\mathcal{D}_{<n-j-1}^{(N-1)}\otimes\mathcal{F}^{n-1})$ (5.6)
$\xrightarrow{\sim}(\mathcal{D}_{<n-j}^{(N-1)}/\mathcal{D}_{<n-j-1}^{(N-1)})\otimes\mathcal{F}^{n-1}$
$\xrightarrow{\sim}\mathcal{T}^{\otimes(n-j-1)}\otimes\mathcal{F}^{n-1}$
for every $j=0,\cdots,n-1$. Hence, we obtain a composite of canonical isomorphisms
$\mathrm{det}(\mathcal{F})\xrightarrow{\sim}\bigotimes_{j=0}^{n-1}\mathcal{F}^{j}/\mathcal{F}^{j+1}\xrightarrow{\sim}\bigotimes_{j=0}^{n-1}\mathcal{T}^{\otimes(n-j-1)}\otimes\mathcal{F}^{n-1}\xrightarrow{\sim}\mathcal{T}^{\otimes\frac{n(n-1)}{2}}\otimes(\mathcal{F}^{n-1})^{\otimes n}.$ (5.7)
In particular, the relative degree $\mathrm{deg}(\mathcal{L})$ of $\mathcal{L}$ can be calculated as follows:
$\mathrm{deg}(\mathcal{L})$ $=\mathrm{deg}(\mathcal{T}^{\otimes\frac{n(n-1)}{2}}\otimes(\mathcal{F}^{n-1})^{\otimes n})$ (5.8)
$=-\frac{n(n-1)(2g-2+r)}{2}+n\cdot\mathrm{deg}(\mathcal{F}^{n-1}).$
It follows that $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ becomes empty when $n$ does not divide the integer $\mathrm{deg}(\mathcal{L})+\frac{n(n-1)(2g-2+r)}{2}$.
5.2. Relation with dormant $\mathrm{PGL}_{n}^{(N)}$-opers
This subsection describes the relationship between $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ and the moduli stack of dormant $\mathrm{PGL}_{n}^{(N)}$-opers. The precise definition of a dormant $\mathrm{PGL}_{n}^{(N)}$-oper (of prescribed radii) is omitted here, so we refer the reader to  [Wak4, § 5] for details. Note that we will only use the “N=1N=1” case in the subsequent discussion (cf. Theorem 6.7, (i)), and that case is comprehensively investigated in  [Wak3].
Hereinafter, we assume that $n<p$. We shall denote by
$\overline{\Xi}_{n,N}^{<}$ (5.9)
the set of equivalence classes of elements in $\Xi_{n,N}^{<}$, where the equivalence relation is defined as follows: $(a^{[1]},\cdots,a^{[n]})$ and $(b^{[1]},\cdots,b^{[n]})$ are equivalent if there exists an element $c\in\mathbb{Z}$ such that the two subsets of $\mathbb{Z}/p^{N}\mathbb{Z}$ defined, via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$, by the sets $\{a^{[j]}+c\}_{j=1}^{n}$ and $\{b^{[j]}\}_{j=1}^{n}$ coincide. (Note that $\overline{\Xi}_{n,N}^{<}$ may be regarded as a subset of $\Xi_{n,N}$” introduced in  [Wak4, § 6.3].) This set has a natural projection
$\rho:\Xi_{n,N}^{<}\twoheadrightarrow\overline{\Xi}_{n,N}^{<}.$ (5.10)
The assumption $n<p$ implies that, for each $i\in\{1,\cdots,r\}$ and $d:=(d^{[1]},\cdots,d^{[n]})\in\overline{\Xi}_{n,N}^{<}$, there exists a unique $(b^{[1]},\cdots,b^{[n]})\in\rho^{-1}(d)$ such that $\sum\limits_{j=1}^{n}b^{[j]}$ coincides with the exponent of $\mathcal{L}_{\flat}$ at $\sigma_{i}$.
Let $\vec{a}:=(a_{1},\cdots,a_{r})$ be as before, and write $\rho(\vec{a}):=(\rho(a_{1}),\cdots,\rho(a_{r}))$. According to  [Wak4, § 6.3], one can obtain the moduli stack
$\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}$ (5.11)
classifying dormant $\mathrm{PGL}_{n}^{(N)}$-opers on $\mathscr{X}$ of radii $\rho(\vec{a})$. (That is to say, $\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}$ is the fiber product of the projection $\Pi_{\rho(\vec{a}),g,r}:\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\rho(\vec{a}),g,r}\rightarrow\overline{\mathcal{M}}_{g,r}$” introduced in  [Wak4, § 6.3] and the classifying morhism $S\rightarrow\mathcal{M}_{g,r}\subseteq\overline{\mathcal{M}}_{g,r}$ of $\mathscr{X}$. Here, $\mathcal{M}_{g,r}$ denotes the moduli stack of $r$-pointed smooth proper curves of genus $g$ over $k$, and $\overline{\mathcal{M}}_{g,r}$ denotes its Deligne-Mumford compactification). Each dormant $\mathrm{GL}_{n}^{(N)}$-oper induces a dormant $\mathrm{PGL}_{n}^{(N)}$-oper via projectivization, and this assignment yields a morphism of $S$-stacks
$\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}\rightarrow\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}.$ (5.12)
Next, for each $d\in\mathbb{Q}$, denote by $\mathcal{P}ic^{d}_{X/S}$ the relative Picard stack of $X/S$ classifying line bundles on $X$ of relative degree $d$, where $\mathcal{P}ic^{d}_{X/S}:=\emptyset$ if $d\in\mathbb{Q}\setminus\mathbb{Z}$. We set $q:=\mathrm{deg}(\mathcal{L})+\frac{n(n-1)(2g-2+r)}{2}$. Then, the assignment $\mathcal{N}\mapsto\mathcal{N}^{\otimes n}$ determines a morphism of $S$-stacks
$[n]:\mathcal{P}ic_{X/S}^{q/n}\rightarrow\mathcal{P}ic_{X/S}^{q}.$ (5.13)
The fiber of this morphism over the $S$-rational point classifying $\mathcal{L}\otimes\Omega^{\otimes\frac{n(n-1)}{2}}$ defines a (possibly empty) $S$-stack
$\mathcal{R}_{n,\mathcal{L}}.$ (5.14)
The structure morphism $\mathcal{R}_{n,\mathcal{L}}\rightarrow S$ of $\mathcal{R}_{n,\mathcal{L}}$ is quasi-finite, proper, and étale (by the assumption $n<p$). Also, it follows from (5.8) that the assignment $(\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j}),\eta)\mapsto\mathcal{F}^{n-1}$ gives an $S$-morphism
$\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}\rightarrow\mathcal{R}_{n,\mathcal{L}}.$ (5.15)
Proposition 5.3.
(Recall that we have assumed the inequality $n<p$.) The following assertions hold.
  • (i)
    The morphism
[…]
    is an isomorphism of $S$-stacks.
  • (ii)
    The structure morphism $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}\rightarrow S$ of $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ is (not representable but) quasi-finite and proper. Moreover, if $n=2$, then this morphism is faithfully flat.
Proof.
To prove assertion (i), we construct an inverse morphism of (5.16). The problem is to construct a functor
$\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}(S)\times\mathcal{R}_{n,\mathcal{L}}(S)\rightarrow\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}(S)$ (5.17)
[…]
Let us take a pair $(\mathscr{E}^{\spadesuit},\mathcal{Q})$ consisting of a dormant $\mathrm{PGL}_{n}^{(N)}$-oper $\mathscr{E}^{\spadesuit}$ on $\mathscr{X}$ of radii $\rho(\vec{a})$ and a line bundle $\mathcal{Q}$ on $X$ classified by $\mathcal{R}_{n,\mathcal{L}}$. By definition, $\mathcal{Q}$ admits an isomorphism $\eta:\mathcal{T}^{\otimes\frac{n(n-1)}{2}}\otimes\mathcal{Q}^{\otimes n}\xrightarrow{\sim}\mathcal{L}$. Then, $\nabla_{\mathcal{L}}$ corresponds, via this isomorphism, to a $\mathcal{D}^{(N-1)}$-module structure $\nabla_{\vartheta}$ on $\mathcal{T}^{\otimes\frac{n(n-1)}{2}}\otimes\mathcal{Q}^{\otimes n}$. The pair $\vartheta:=(\mathcal{Q},\nabla_{\vartheta})$ specifies a (dormant) $n^{(N)}$-theta characteristic of $\mathscr{X}$, in the sense of  [Wak4, § 5.4]. By  [Wak4, Theorem 5.5.1], there exists a unique dormant $(\mathrm{GL}_{n}^{(N)},\vartheta)$-oper inducing $\mathscr{E}^{\spadesuit}$ via projectivization; in particular, this defines a dormant $\mathrm{GL}^{(N)}_{n}-$oper $\mathscr{F}^{\heartsuit}$, and $\eta$ forms an isomorphism between its determinant and $\mathcal{L}_{\flat}$. Thus, the pair $(\mathscr{F}^{\heartsuit},\eta)$ specifies an object of $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}(S)$ (cf. the italicized comment in the second paragraph of this subsection). The resulting assignment $(\mathscr{E}^{\spadesuit},\mathcal{Q})\mapsto\mathscr{F}^{\heartsuit}$ turns out to define the desired functor (5.17). This completes the proof of assertion (i).
Moreover, assertion (ii) follows from  [Wak4, Theorems B and C]. ∎
Corollary 5.4.
(Recall that we have assumed the inequality $n<p$.) Suppose the following two conditions:
  • The integer $\mathrm{deg}(\mathcal{L})+\frac{n(n-1)(2g-2+r)}{2}$ is divided by $n$;
  • $S=\mathrm{Spec}(k)$, i.e., $\mathscr{X}$ is an $r$-pointed smooth proper curve over $k$ of genus $g$.
Then, the following equality holds:
$\sharp(\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}(k))=n^{2g}\cdot\sharp(\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}(k))<\infty,$ (5.18)
where for a category $\mathcal{C}$ we denote by $\sharp(\mathcal{C})$ the cardinality of the set of isomorphism classes of objects in $\mathcal{C}$.
Proof.
Since the first assumption implies $\sharp(\mathcal{R}_{n,\mathcal{L}}(k))=n^{2g}$, the assertion follows from Proposition 5.3, (i). ∎
Remark 5.5.
According to an argument similar to  [Wak3, Remark 8.21] together with Proposition 5.3, (i) and (ii), there exists an open substack
${{}^{\circledcirc}}\mathcal{M}_{g,r}$ (5.19)
of $\mathcal{M}_{g,r}$ classifying pointed curves $\mathscr{X}$ such that $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}}$ is étale whenever $\mathcal{L}_{\flat}$ is a $p^{N}$-flat line bundle whose exponent at $\sigma_{i}$ coincides with $\sum\limits_{j=1}^{n}a_{i}^{[j]}$ modulo $p^{N}$ for every $i$. Since $\mathcal{M}_{g,r}$ is irreducible, ${{}^{\circledcirc}}\mathcal{M}_{g,r}$ is either empty or dense in $\mathcal{M}_{g,r}$. In the case of $n=2$, we can find at least one choice of $\vec{a}$ such that ${{}^{\circledcirc}}\mathcal{M}_{g,r}$ is dense (cf.  [Wak4, Theorems B and C]).
5.3. Harder-Narasimhan polygons
We shall consider a characterization of $\mathrm{GL}_{n}^{(N)}$-opers by using their Harder-Narasimhan polygons; our result is a higher-level and parabolic generalization of  [JoPa, Theorem 1.1.2]. In this subsection, we suppose that $S=\mathrm{Spec}(k)$, i.e., $\mathscr{X}$ is an $r$-pointed smooth proper curve of genus $g$ over $k$.
Recall that each vector bundle $\mathcal{F}$ on $X$ admits the Harder-Narasimhan filtration
$0=\mathcal{F}^{m}\subsetneq\mathcal{F}^{m-1}\subsetneq\cdots\subsetneq\mathcal{F}^{1}\subsetneq\mathcal{F}^{0}=\mathcal{F}.$ (5.20)
That is to say, $\mathcal{F}^{j}/\mathcal{F}^{j+1}$ are all semistable, and the slopes (in the usual sense) of these vector bundles satisfy
$\mu_{\mathrm{max}}(\mathcal{F}):=\mu(\mathcal{F}^{m-1}/\mathcal{F}^{m})>\mu(\mathcal{F}^{m-2}/\mathcal{F}^{m-1})>\cdots>\mu(\mathcal{F}^{0}/\mathcal{F}^{1})=:\mu_{\mathrm{min}}(\mathcal{F}).$ (5.21)
This filtration associates a polygon in the plane $\mathbb{R}^{2}$ whose vertices are given by the points $(\mathrm{rk}(\mathcal{F}^{j}),\mathrm{deg}(\mathcal{F}^{j}))$ ($0\leq j\leq m$). We denote this polygon by
$P_{\mathcal{F}}.$ (5.22)
The segment of $P_{\mathcal{F}}$ joining $(\mathrm{rk}(\mathcal{F}^{j}),\mathrm{deg}(\mathcal{F}^{j}))$ and $(\mathrm{rk}(\mathcal{F}^{j-1}),\mathrm{deg}(\mathcal{F}^{j-1}))$ has slope $\mu(\mathcal{F}^{j-1}/\mathcal{F}^{j})$.
Let us consider the set $B$ of convex polygons in $\mathbb{R}^{2}$ starting at $(0,0)$ and ending at $(n,na-\frac{n(n-1)(2g-2+r)}{2})$; this set is partially ordered: if $P_{1}$, $P_{2}$ are two such polygons, we say that $P_{1}\succcurlyeq P_{2}$ if $P_{1}$ lies on or above $P_{2}$. Also, this set contains a specific polygon with vertices $q_{j}:=\left(j,ja-\frac{j(j-1)(2g-2+r)}{2}\right)$ for $0\leq j\leq n$; we denote it by
$P_{\mathrm{oper}}^{a}.$ (5.23)
As asserted in the proposition below, this convex polygon characterizes $\mathcal{D}^{(N-1)}$-modules admitting a structure of $\mathrm{GL}_{n}^{(N)}$-oper.
Proposition 5.6 (cf.  [JoPa], Theorem 1.1.2).
Let $(m_{1},\cdots,m_{r})$ be an $r$-tuple of positive integers and $\vec{a}:=(a_{1},\cdots,a_{r})$ an element of $\prod_{i=1}^{r}\Xi_{m_{i},N}^{\leq}$ such that $s_{1}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},1}^{\leq}$ and $s_{2}^{1}(\vec{a})\in\prod_{i=1}^{r}\Xi_{m_{i},N-1}^{\leq}$ (cf. (4.25)). Also, let $n$ be a positive integer and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ a rank-$n$ parabolic bundle on $\mathscr{X}^{(N)}$ of parabolic weights $\vec{a}/p^{N}$ whose parabolic Frobenius pull-back by $F^{(N-1)}_{X/S}$ is semistable. Denote by $\mathscr{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ the parabolic Frobenius pull-back of $\mathscr{E}$ by $F^{(N)}_{X/S}$ and by
$0=\mathcal{F}^{m}\subsetneq\mathcal{F}^{m-1}\subsetneq\cdots\subsetneq\mathcal{F}^{0}=\mathcal{F}$ (5.24)
the Harder-Narasimhan filtration of $\mathcal{F}:=\mathcal{E}^{F}$. Suppose that $\mathcal{F}$ has degree $na-\frac{n(n-1)(2g-2+r)}{2}$ for some integer $a$. Then, the following assertions hold.
  • (i)
    We have $P_{\mathrm{oper}}^{a}\succcurlyeq P_{\mathcal{F}}$.
  • (ii)
    The equality $P_{\mathrm{oper}}^{a}=P_{\mathcal{F}}$ holds if and only if the triple $(\mathcal{F},\nabla^{F},\{\mathcal{F}^{j}\}_{j})$ forms a (dormant) $\mathrm{GL}_{n}^{(N)}$-oper. (If these conditions are fulfilled, then $n=m$ and the equality $\mathrm{deg}(\mathcal{F}^{n-1})=a$ holds.)
Proof.
First, we shall prove assertion (i). Suppose that there exists a vertex of $P_{\mathcal{F}}$ that lies above $P_{\mathrm{oper}}^{a}$. The segment $S_{j}$ joining the vertices $q_{j-1}$ and $q_{j}$ has slope $s_{j}:=a-(j-1)(2g-2+r)$. Since $s_{j}-s_{j+1}=2g-2+r$, it follows from an elementary argument about slopes of convex polygons that there exists $j\in\{1,\cdots,m-1\}$ satisfying $\mu(\mathcal{F}^{j}/\mathcal{F}^{j+1})-\mu(\mathcal{F}^{j-1}/\mathcal{F}^{j})>2g-2+r$. This contradicts the result of Lemma 5.7 proved below. Thus, we have $P_{\mathrm{oper}}^{a}\succcurlyeq P_{\mathcal{F}}$, as desired.
Next, we shall prove assertion (ii). Since the “if” part of the required equivalence is clear, we only consider the “only if” part. Suppose that $P_{\mathrm{oper}}^{a}=P_{\mathcal{F}}$ (hence $m=n$). As an intermediate step, we shall prove, by descending induction on $j$, the claim that $\nabla^{F}_{0}(\mathcal{F}^{j})\subseteq\mathcal{F}^{j-1}$ ($j=1,\cdots,n$), where $\nabla^{F}_{0}$ denotes the log connection on $\mathcal{F}$ induced from $\nabla^{F}$.
The base step, i.e., the case of $j=n$, is clear. To consider the induction step, we suppose that the claim with $j$ replaced by $j+1$ has been proved. The composite
$\alpha:\mathcal{F}^{j}\xrightarrow{\mathrm{inclusion}}\mathcal{F}\xrightarrow{\nabla^{F}_{0}}\Omega\otimes\mathcal{F}\xrightarrow{\mathrm{quotient}}\Omega\otimes(\mathcal{F}/\mathcal{F}^{j})$ (5.25)
is verified to be $\mathcal{O}_{X}$-linear. By the induction assumption, $\alpha$ induces, via the quotient $\mathcal{F}^{j}\twoheadrightarrow\mathcal{F}^{j}/\mathcal{F}^{j+1}$, an $\mathcal{O}_{X}$-linear morphism $\overline{\alpha}:\mathcal{F}^{j}/\mathcal{F}^{j+1}\rightarrow\Omega\otimes(\mathcal{F}/\mathcal{F}^{j})$. Since $\alpha$ (hence also $\overline{\alpha}$) is nonzero by Lemma 5.7, $\mathrm{Im}(\overline{\alpha})$ turns out to be a rank one vector bundle. Thus, there exists at least one integer $s<j$ such that the image of $\mathrm{Im}(\overline{\alpha})$ via the natural quotient $\Omega\otimes(\mathcal{F}/\mathcal{F}^{j})\twoheadrightarrow\Omega\otimes(\mathcal{F}/\mathcal{F}^{s+1})$ is nonzero; let $s_{\mathrm{min}}$ be the minimum element among such $s$’s. By definition, $\overline{\alpha}$ induces an injection $\mathcal{F}^{j}/\mathcal{F}^{j+1}\hookrightarrow\Omega\otimes(\mathcal{F}^{s_{\mathrm{min}}}/\mathcal{F}^{s_{\mathrm{min}}+1})$. But, the following equalities hold:
$\ \ \ \ \mathrm{deg}(\Omega\otimes(\mathcal{F}^{s_{\mathrm{min}}}/\mathcal{F}^{s_{\mathrm{min}}+1}))-\mathrm{deg}(\mathcal{F}^{j}/\mathcal{F}^{j+1})$ (5.26)
$=\left((2g-2+r)+a-(n-s_{\mathrm{min}}-1)(2g-2+r)\right)-\left(a-(n-j-1)(2g-2+r)\right)$
$=(2g-2+r)(s_{\mathrm{min}}-j+1).$ (5.27)
It follows that the equality $s_{\mathrm{min}}=j-1$ must be fulfilled, which means $\nabla^{F}_{0}(\mathcal{F}^{j})\subseteq\mathcal{F}^{j-1}$. This completes the proof of the induction step, and hence completes the proof of the claim.
By comparing the respective degrees, we see that the injection $\mathcal{F}^{j}/\mathcal{F}^{j+1}\xrightarrow{}\Omega\otimes(\mathcal{F}^{j-1}/\mathcal{F}^{j})$ induced by $\nabla^{F}_{0}$ is bijective. Therefore, the morphism (5.3) for $(\mathcal{F},\nabla^{F},\{\mathcal{F}^{j}\}_{j})$ turns out to be an isomorphism. That is to say, it forms a $\mathrm{GL}_{n}^{(N)}$-oper. This completes the proof of the “only if” part. ∎
The following lemma was used in the proof of the above proposition (cf.  [JoPa, Lemma 5.1.1],  [LasPa, Lemma 4.2],  [She, Corollary 2],  [Sun, Theorem 3.1]).
Lemma 5.7.
Let us keep the notation and assumption in the above proposition. Then, for each $j=1,\cdots,m-1$, the subbundle $\mathcal{F}^{j}$ of $\mathcal{F}$ is not closed under $\nabla^{F}$ and satisfies the inequality
$\mu(\mathcal{F}^{j}/\mathcal{F}^{j+1})-\mu(\mathcal{F}^{j-1}/\mathcal{F}^{j})\leq 2g-2+r.$ (5.28)
In particular, (since $m<\mathrm{rank}(\mathcal{E})$) we have
$\mu_{\mathrm{max}}(\mathcal{F})-\mu_{\mathrm{min}}(\mathcal{F})\leq(\mathrm{rank}(\mathcal{E})-1)(2g-2+r).$ (5.29)
Proof.
By the assumption on $\vec{a}$ together with Proposition 4.7, we may assume that $N=1$. Let us choose $j\in\{1,\cdots,m-1\}$. If $\nabla_{0}^{F}$ is as in the proof of the above proposition, then the composite
$\mathcal{F}^{j}\xrightarrow{\mathrm{inclusion}}\mathcal{F}\xrightarrow{\nabla^{F}_{0}}\Omega\otimes\mathcal{F}\xrightarrow{\mathrm{quotient}}\Omega\otimes(\mathcal{F}/\mathcal{F}^{j})$ (5.30)
is verified to be $\mathcal{O}_{X}$-linear.
Suppose that $\mathcal{F}^{j}$ is closed under $\nabla^{F}_{0}$, i.e., (5.30) becomes the zero map. It follows from Proposition 4.6 that the parabolic $p$-flat subbundle $\mathscr{E}^{F,j}$ determined by $\mathcal{F}^{j}$ corresponds to a parabolic subbundle $\mathscr{G}$ of $\mathscr{E}$. By (4.22) and the definition of parabolic slope, we have
$\mathrm{par}\text{-}\mu(\mathscr{E})$ $=\frac{1}{p}\cdot\left(\mathrm{par}\text{-}\mu(\mathscr{E}^{F})\right)$ (5.31)
$=\frac{1}{p}\cdot\left(\mu(\mathcal{F})-\frac{1}{n}\left(\sum_{i=1}^{r}\sum_{j^{\prime}=1}^{m_{i}}a_{i}^{[j^{\prime}]}\ell_{i}^{[j^{\prime}]}\right)\right)$
$<\frac{1}{p}\cdot\left(\mu(\mathcal{F}^{j})-\frac{1}{n}\cdot\left(\sum_{i=1}^{r}\sum_{j^{\prime}=1}^{m_{i}}a_{i}^{[j^{\prime}]}\ell_{i,\mathcal{F}^{j}}^{[j^{\prime}]}\right)\right)$
$=\frac{1}{p}\cdot\mathrm{par}\text{-}\mu(\mathscr{E}^{F,j})$
$=\mathrm{par}\text{-}\mu(\mathscr{G})$
where the “<<” follows from the fact that $\{\mathcal{F}^{j}\}_{j}$ is the Harder-Narasimhan filtration (which implies $\mu(\mathcal{F}^{j})>\mu(\mathcal{F})$). This contradicts the semistability assumption on $\mathscr{E}$, and hence, proves the first assertion.
Moreover, this fact means that (5.30) is nonzero, which implies
$\mu(\mathcal{F}^{j}/\mathcal{F}^{j+1})=\mu_{\mathrm{min}}(\mathcal{F}^{j})\leq\mu_{\mathrm{max}}(\Omega\otimes(\mathcal{F}/\mathcal{F}^{j}))=\mu(\mathcal{F}^{j-1}/\mathcal{F}^{j})+2g-2+r.$ (5.32)
This completes the proof of the second assertion. ∎
6. Maximally Frobenius-destabilized bundles
Maximally Frobenius-destabilized bundles on unpointed curves were investigated in some references, e.g.,  [JoPa],  [JRXY],  [JoXi],  [LanPa], and  [Zha]. In particular, it has been shown that these bundles correspond bijectively to dormant opers via Cartier descent; this result enables us to resolve related counting problems in the study of vector bundles in positive characteristic. In this section, we describe the higher-level and parabolic generalization of this result (cf. Theorem 6.4).
We suppose that $S=\mathrm{Spec}(k)$, i.e., $\mathscr{X}$ is an $r$-pointed smooth proper curve of genus $g$ over $k$.
6.1. Maximally $F^{(N)}$-destabilized bundles
Let $n$ be a positive integer and $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ (where $\vec{\mathfrak{e}}:=(\mathfrak{e}_{1},\cdots\mathfrak{e}_{r})$) a rank-$n$ parabolic bundle on $\mathscr{X}^{(N)}$ such that the type of the parabolic structure $\mathfrak{e}_{i}$ at $\sigma_{i}$ coincides with $1_{\times n}:=(1,\cdots,1)\in\mathbb{Z}_{>0}^{\times n}$ for every $i=1,\cdots,r$ and $\vec{a}$ belongs to $(\Xi_{n,N}^{\leq})^{\times r}$. This bundle gives rise to its parabolic Frobenius pull-back $\mathscr{E}^{F}:=(\mathcal{E}^{F},\nabla^{F},\vec{\mathfrak{e}}^{F},\vec{a})$ by $F^{(N)}_{X/S}$. For convenience, we write $\mathcal{F}:=\mathcal{E}^{F}$.
Definition 6.1.
We shall say that $\mathscr{E}$ is maximally $F^{(N)}$-destabilized if there exists an $n$-step decreasing filtration
$0=\mathcal{F}^{n}\subsetneq\mathcal{F}^{n-1}\subsetneq\cdots\subsetneq\mathcal{F}^{1}\subsetneq\mathcal{F}^{0}=\mathcal{F}$ (6.1)
such that (the subquotients $\mathcal{F}^{j}/\mathcal{F}^{j+1}$ are line bundles and) the equality
$\left(\mu(\mathcal{F}^{j}/\mathcal{F}^{j+1})-\mu(\mathcal{F}^{j-1}/\mathcal{F}^{j})=\right)\mathrm{deg}(\mathcal{F}^{j}/\mathcal{F}^{j+1})-\mathrm{deg}(\mathcal{F}^{j-1}/\mathcal{F}^{j})=2g-2+r$ (6.2)
holds for every $j=1,\cdots,m-1$. We shall refer to such a filtration as a destabilizing filtration of $\mathscr{E}$.
Proposition 6.2.
Let $\mathscr{E}$ be as above, and let us consider the following three conditions:
  • (a)
    $\mathscr{E}$ is maximally $F^{(N)}$-destabilized;
  • (b)
    $s_{1}^{1}(\vec{a})\in(\Xi_{n,1}^{\leq})^{\times r}$ and $s_{2}^{1}(\vec{a})\in(\Xi_{n,N-1}^{\leq})^{\times r}$ (cf. (4.25));
  • (c)
    The inequalities $\sum\limits_{i=1}^{r}(s_{1}^{1}(a_{i}^{[n]})-s_{1}^{1}(a_{i}^{[1]}))<\frac{n(2g-2+r)}{2}<\frac{p}{n}$ hold.
Then, the following assertions hold.
  • (i)
    If (a) is satisfied, then a destabilizing filtration of $\mathscr{E}$ is uniquely determined. (Thus, it makes sense to speak of “the” destabilizing filtration of $\mathscr{E}$.)
  • (ii)
    If $n\mid\mathrm{deg}(\mathcal{E})$ and all the conditions (a)-(c) are satisfied, then $\mathscr{E}$ is stable.
Proof.
Assertion (i) can be proved by a routine argument using 2g2+r>02g-2+r>0 (so we omit the detail of the proof).
In the following, we prove assertion (ii). Suppose that there exists a proper parabolic subbundle $\mathscr{G}:=(\mathcal{G},\vec{\mathfrak{g}},\vec{a}/p^{N})$ of $\mathscr{E}$ satisfying $\mathrm{par}\text{-}\mu(\mathscr{G})\geq\mathrm{par}\text{-}\mu(\mathscr{E})$. By Proposition 4.6, the parabolic Frobenius pull-back $\mathscr{G}^{F}:=(\mathcal{G}^{F},\nabla^{F}_{\mathcal{G}},\vec{\mathfrak{g}}^{F},\vec{a})$ specifies a parabolic $p^{N}$-flat subbundle of $\mathscr{E}^{F}$. Denote by $m\left(<n\right)$ the rank of $\mathcal{G}^{F}$, and define a decreasing filtration $\{\mathcal{G}^{F,j}\}_{j=0}^{n}$ on $\mathcal{G}^{F}$ by putting $\mathcal{G}^{F,j}:=\iota^{-1}(\mathcal{F}^{j})$, where $\{\mathcal{F}^{j}\}_{j=0}^{n}$ denotes the destabilized filtration of $\mathscr{E}$ and $\iota$ denotes the natural inclusion $\mathcal{G}^{F}\hookrightarrow\mathcal{F}\left(=\mathcal{F}^{0}=\mathcal{E}^{F}\right)$. Then, one obtains a subset $B:=\left\{j\,|\,\mathcal{G}^{F,j}\neq\mathcal{G}^{F,j+1}\right\}$ of $\{0,\cdots,n-1\}$. By Lemma 6.3 below together with an argument similar to the proof of Proposition 5.6 (by using the assumption (b)), we see that the log connection on $\mathcal{F}$ induced from $\nabla^{F}$ gives, in a usual manner, an injection $\mathcal{F}^{j}/\mathcal{F}^{j+1}\hookrightarrow\Omega\otimes(\mathcal{F}^{j-1}/\mathcal{F}^{j})$ (i.e., the $j$-th Kodaira-Spencer map). By the condition (a), this morphism is verified to be an isomorphism; it moreover restricts to an injection $\mathcal{G}^{F,j}/\mathcal{G}^{F,j+1}\hookrightarrow\Omega\otimes(\mathcal{G}^{F,j-1}/\mathcal{G}^{F,j})$. It follows that $j\in B\cap\{1,\cdots,n-1\}$ implies $j-1\in B$, and hence that $B=\{0,\cdots,m-1\}$. Hence, the composite $\mathcal{G}^{F}\xrightarrow{\mathrm{inclusion}}\mathcal{F}\twoheadrightarrow\mathcal{F}/\mathcal{F}^{m}$ turns out to be injective. This implies
$\mu(\mathcal{G}^{F})\leq\mu(\mathcal{F}/\mathcal{F}^{m}).$ (6.3)
On the other hand, if we write $c:=\mathrm{deg}(\mathcal{F}^{n-1})$, then we have
$\mathrm{par}\text{-}\mu(\mathscr{E})$ $=\frac{1}{p^{N}\cdot n}\cdot\mathrm{deg}(\mathcal{F})$ (6.4)
$=\frac{1}{p^{N}\cdot n}\cdot\sum_{j=0}^{n-1}\mathrm{deg}(\mathcal{F}^{j}/\mathcal{F}^{j+1})$
$=\frac{1}{p^{N}\cdot n}\cdot\sum_{j=0}^{n-1}\left(c-\frac{(n-j-1)(2g-2+r)}{2}\right)$
$=\frac{1}{p^{N}}\cdot\left(c-\frac{(n-1)(2g-2+r)}{2}\right),$
where the first equality follows from (4.22) and the third equality follows from the assumption that $\mathscr{E}$ is maximally $F^{(N)}$-destabilized. Thus, by combining it with (6.3), we have
$\mathrm{par}\text{-}\mu(\mathscr{G})$ $=\frac{1}{p^{N}}\cdot\mu(\mathcal{G}^{F})$ (6.5)
$\leq\frac{1}{p^{N}}\cdot\mu(\mathcal{F}/\mathcal{F}^{m})$
$=\frac{1}{p^{N}\cdot m}\cdot\sum_{j=0}^{m-1}\mathrm{deg}(\mathcal{F}^{j}/\mathcal{F}^{j+1})$ (6.6)
$=\frac{1}{p^{N}\cdot m}\cdot\sum_{j=0}^{m-1}\left(c-\frac{(n-j-1)(2g-2+r)}{2}\right)$
$=\frac{1}{p^{N}}\cdot\left(c-\frac{(2n-m-1)(2g-2+r)}{2}\right)$
[…]
$<\mathrm{par}\text{-}\mu(\mathscr{E}).$ (6.7)
This is a contradiction, and we finish the proof of assertion (ii). ∎
The following lemma was used in the proof of the above proposition.
Lemma 6.3.
Suppose that $n\mid\mathrm{deg}(\mathcal{E})$ and the conditions (a)-(c) in the above proposition are satisfied. Then, the proper nontrivial subbundles $\mathcal{F}^{j}$ ($j=1,\cdots,n-1$) of $\mathcal{F}$ are not closed under $\nabla^{F}$.
Proof.
To prove the assertion, we are always free to replace $S$ with its étale covering. Hence, since $n\mid\mathrm{deg}(\mathcal{E})$, we may assume, after possibly tensoring $\mathcal{E}$ with a line bundle, that $\mathrm{deg}(\mathcal{E})=0$. By the condition (b) together with Proposition 4.7, we may assume that $N=1$.
[…]
$\mathrm{deg}(\mathcal{F})$ $=\sum_{j^{\prime}=0}^{n-1}\mathrm{deg}(\mathcal{F}^{j^{\prime}}/\mathcal{F}^{j^{\prime}+1})$ (6.8)
$=\sum_{j^{\prime}=0}^{n-1}\left(\mathrm{deg}(\mathcal{F}^{n-1})-(n-j^{\prime}-1)(2g-2+r)\right)$ (6.9)
$=n\cdot\mathrm{deg}(\mathcal{F}^{n-1})-\frac{n(n-1)(2g-2+r)}{2}.$
Hence, the equality $\mathrm{deg}(\mathcal{E})=0$ and (4.21) together imply
$\sum_{i=1}^{r}\sum_{j=1}^{n}a_{i}^{[j]}\left(=p\cdot\mathrm{deg}(\mathcal{E})+\sum_{i=1}^{r}\sum_{j=1}^{n}a_{i}^{[j]}\right)=n\cdot\mathrm{deg}(\mathcal{F}^{n-1})-\frac{n(n-1)(2g-2+r)}{2}.$ (6.10)
It follows that
$\mathrm{deg}(\mathcal{F}^{j})$ $=\sum_{j^{\prime}=j}^{n-1}\mathrm{deg}(\mathcal{F}^{j^{\prime}}/\mathcal{F}^{j^{\prime}+1})$ (6.11)
$=\sum_{j^{\prime}=j}^{n-1}\left(\mathrm{deg}(\mathcal{F}^{n-1})-(n-j^{\prime}-1)(2g-2+r)\right)$
$=(n-j)\cdot\mathrm{deg}(\mathcal{F}^{n-1})-\frac{(n-j)(n-j-1)(2g-2+r)}{2}$
$=\frac{j(n-j)(2g-2+r)}{2}+\frac{n-j}{n}\cdot\sum_{i=1}^{r}\sum_{j=1}^{n}a_{i}^{[j]}.$ (6.12)
Next, observe that
$\ \ \ \ \mathrm{deg}(F^{*}_{X/S}\mathcal{S}ol(\nabla)\cap\mathcal{F}^{j})$ (6.13)
$\leq\mathrm{deg}(\mathcal{F}^{j})-\sum_{i=1}^{r}\sum_{j^{\prime}=1}^{n-j}a_{i}^{[j^{\prime}]}$
$=\frac{j(n-j)(2g-2+r)}{2}+\frac{1}{n}\cdot\sum_{i=1}^{r}\left((n-j)\cdot\sum_{j^{\prime}=n-j+1}^{n}a_{i}^{[j]}-j\cdot\sum_{j^{\prime}=1}^{n-j}a_{i}^{[j^{\prime}]}\right)$
$=\frac{1}{n}\cdot\sum_{j^{\prime}=n-j+1}^{n}\sum_{j^{\prime\prime}=1}^{n-j}\left(\frac{n(2g-2+r)}{2}+\sum_{i=1}^{r}(a_{i}^{[j^{\prime}]}-a_{i}^{[j^{\prime\prime}]})\right)$
$<\frac{1}{n}\cdot\frac{n^{2}}{2}\cdot\left(\frac{n(2g-2+r)}{2}+\frac{n(2g-2+r)}{2}\right)$
$\leq p,$ (6.14)
where the inequality “<<” follows from the first inequality in the condition (c) and the last inequality $\leq$” follows from the second one in (c). On the other hand, we have
$\ \ \ \ \mathrm{deg}(F^{*}_{X/S}\mathcal{S}ol(\nabla)\cap\mathcal{F}^{j})$ (6.15)
$\geq\mathrm{deg}(\mathcal{F}^{j})-\sum_{i=1}^{r}\sum_{j^{\prime}=j+1}^{n}a_{i}^{[j^{\prime}]}$
$=\frac{j(n-j)(2g-2+r)}{2}+\frac{1}{n}\cdot\sum_{i=1}^{r}\left((n-j)\cdot\sum_{j^{\prime}=1}^{j}a_{i}^{[j^{\prime}]}-j\cdot\sum_{j^{\prime}=j+1}^{n}a_{i}^{[j^{\prime}]}\right)$
$=\frac{1}{n}\cdot\sum_{j^{\prime}=1}^{j}\sum_{j^{\prime\prime}=j+1}^{n}\left(\frac{n(2g-2+r)}{2}+\sum_{i=1}^{r}(a_{i}^{[j^{\prime}]}-a_{i}^{[j^{\prime\prime}]})\right)$
$>0.$ (6.16)
These computations contradict the fact that $\mathrm{deg}(F^{*}_{X/S}\mathcal{S}ol(\nabla)\cap\mathcal{F}^{j})$ is divided by $p$. This completes the proof of this lemma. ∎
Let $\mathcal{L}$ be a line bundle on $X^{(N)}$ with $n\mid\mathrm{deg}(\mathcal{L})$ and $\vec{a}$ an element of $(\Xi_{n,N}^{<})^{\times r}\left(\subseteq(\Xi_{n,N}^{\leq})^{\times r}\right)$ as before such that the conditions (b) and (c) described in Proposition 6.2 are fulfilled. Also, we set $\vec{1}_{\times n}:=(1_{\times n},\cdots,1_{\times n})$ (= the $r$-tuple of copies of $1_{\times n}$) and $\omega:=(n,\vec{1}_{\times n},\vec{a}/p^{N})$. Then, it follows from Proposition 6.2, (ii), that there exists a fibered (full) subcategory
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$ (6.17)
of $\mathcal{U}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$ classifying maximally $F^{(N)}$-destablized stable parabolic bundles.
6.2. Correspondence with dormant $\mathrm{GL}_{n}^{(N)}$-opers
We keep the notation and assumptions introduced at the end of the previous subsection, and write $\mathcal{L}^{F}_{\flat}:=((F^{(N)*}_{X/S}\mathcal{L})(D_{\star}),\nabla^{\mathrm{can}}_{\mathcal{L},D_{\star}})$ (cf. (3.3)), where $D_{\star}:=\sum\limits_{i=1}^{r}(\sum\limits_{j=1}^{n}a_{i}^{[j]})\sigma_{i}$.
Theorem 6.4 (cf. Theorem B).
  • (i)
    Let $\mathscr{E}:=(\mathcal{E},\vec{\mathfrak{e}},\vec{a}/p^{N})$ be the parabolic bundle classified by an $S$-rational point of $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$. Denote by $\{\mathscr{E}^{F,j}\}_{j=0}^{n}$ be the destabilizing filtration of $\mathscr{E}^{F}$. Then, the collection of data
    $\mathscr{E}^{F,\heartsuit}:=(\mathcal{E}^{F},\nabla^{F},\{\mathcal{E}^{F,j}\}_{j})$ (6.18)
    forms a dormant $\mathrm{GL}_{n}^{(N)}$-oper classified by $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{\alpha},\mathcal{L}^{F}_{\flat}}$ (cf. (5.5)).
  • (ii)
    Suppose that $\sum\limits_{j=1}^{n}a_{i}^{[j]}<p^{N}$ for every $i=1,\cdots,r$. Then, the assignment $\mathscr{E}\mapsto\mathscr{E}^{F,\heartsuit}$ defines an isomorphism of fibered categories
    $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\xrightarrow{\sim}\mathcal{O}p_{\mathscr{X},n,\vec{a},\mathcal{L}^{F}_{\flat}}^{{}^{\mathrm{Zzz...}}}$. (6.19)
  • (iii)
    Let us keep the assumption in (ii). Suppose further that $n$ does not divide the integer $\frac{n(n-1)(2g-2+r)}{2}+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}a_{i}^{[j]}$. Then, $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$ is empty.
Proof.
First, we shall consider assertion (i). By Proposition 6.2, (ii), the parabolic Frobenius pull-back of $\mathscr{E}$ by $F_{X/S}^{(N-1)}$ is stable. Hence, the assertion follows from Proposition 5.6, (ii).
Next, we shall consider assertion (ii). The assignment $\mathscr{E}\mapsto\mathscr{E}^{F,\heartsuit}$ is functorial because of its construction, so it defines a functor
$\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\rightarrow\mathcal{O}p_{\mathscr{X},n,\vec{a},\mathcal{L}^{F}_{\flat}}^{{}^{\mathrm{Zzz...}}}$. (6.20)
We shall construct its inverse. Let $(\mathscr{F}^{\heartsuit},\eta)$ (where $\mathscr{F}^{\heartsuit}:=(\mathcal{F},\nabla,\{\mathcal{F}^{j}\}_{j=0}^{n})$) be an object classified by $\mathcal{O}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\vec{a},\mathcal{L}_{\flat}^{F}}$. Apply Lemma 6.5 described below to the restriction of $(\mathcal{F},\nabla)$ to the formal neighborhood of each $\sigma_{i}$. Then, there exists a quasi-parabolic structure $\vec{\mathfrak{f}}$ on $\mathcal{F}$, for which $(\mathcal{F},\nabla,\vec{\mathfrak{f}},\vec{a})$ specifies a parabolic $p^{N}$-flat bundle. It follows from Theorem 4.5 that the horizontal parabolic bundle $\mathscr{F}^{\nabla}:=(\mathcal{F}^{\nabla},\vec{\mathfrak{f}}^{\nabla},\vec{a}/p^{N})$ associated to this collection is maximally $F^{(N)}$-destabilized. Moreover, by the assumptions on $a_{i}^{[j]}$’s imposed in (ii), we see that $\mathscr{F}^{\nabla}$ belongs to $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}(S)$ (cf. Proposition 6.2, (ii)). The formation of the assignment $(\mathscr{F}^{\heartsuit},\eta)\mapsto\mathscr{F}^{\nabla}$ is functorial with respect to $S$, and hence, determines a functor
$\mathcal{O}p_{\mathscr{X},n,\vec{a},\mathcal{L}^{F}_{\flat}}^{{}^{\mathrm{Zzz...}}}\rightarrow\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}.$ (6.21)
By Theorem 4.5, this functor defines an inverse to (6.20). So we finish the proof of assertion (ii).
Finally, since $\frac{n(n-1)(2g-2+r)}{2}+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}a_{i}^{[j]}$ equals the degree of $(F_{X/S}^{(N)*}\mathcal{L})(D_{\star})$ modulo $n$, assertion (iii) follows from assertion (ii) and the fact mentioned at the end of Remark 5.2
The following lemma was used in the proof of the above theorem.
Lemma 6.5.
Let $(\mathcal{F},\nabla)$ be a rank-$n$ $p^{N}$-flat bundle on $\mathscr{U}$ whose exponent is given by $a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]}]$ for some element $a_{\circ}:=(a_{\circ}^{[1]},\cdots,a_{\circ}^{[n]})$ of $\Xi_{n,N}^{<}$. Then, there exists a unique quasi-parabolic structure $\mathfrak{f}_{\circ}$ on $\mathcal{F}$ for which $(\mathcal{F},\nabla,\mathfrak{f}_{\circ})$ forms a parabolic $p^{N}$-flat bundles.
Proof.
Denote by $d^{[j]}$ ($j=1,\cdots,n$) the image of $a_{\circ}^{[j]}$ via the quotient $\mathbb{Z}\twoheadrightarrow\mathbb{Z}/p^{N}\mathbb{Z}$. To prove this assertion, we may assume, without loss of generality, that there exists an isomorphism of $\mathcal{D}_{\circ}^{(N-1)}$-modules $\xi:(\mathcal{F},\nabla)\xrightarrow{\sim}\bigoplus_{j=1}^{n}\mathscr{O}_{d^{[j]},\flat}$. By considering the eigenspaces of the monodromy operator of $\bigoplus_{j=1}^{n}\nabla_{d^{[j]}}$ at $\sigma_{\circ}$, it is verified that there exists a unique quasi-parabolic structure $\mathfrak{f}^{\prime}_{\circ}$ on $\mathcal{O}_{U}^{\oplus n}$ for which the collection $(\mathcal{O}_{U}^{\oplus n},\bigoplus_{j=1}^{n}\nabla_{d^{[j]}},\mathfrak{f}^{\prime}_{\circ})$ forms a parabolic $p^{N}$-flat bundle. The quasi-parabolic structure $\mathfrak{f}_{\circ}$ on $\mathcal{F}$ corresponding to $\mathfrak{f}^{\prime}_{\circ}$ via $\xi$ provides a required one.
Moreover, since $a_{\circ}^{[1]}<\cdots<a_{\circ}^{[n]}$, it follows from the definition of $\mathfrak{f}^{\prime}_{\circ}$ and  [Wak4, Proposition 4.3.4, (ii)] that $\mathfrak{f}^{\prime}_{\circ}$ is invariant under transposing by any automorphism of $\bigoplus_{i=1}^{n}\mathscr{O}_{d_{\circ}^{[j]},\flat}$. Hence, $\mathfrak{f}_{\circ}$ does not depend on the choice of $\xi$. This proves the uniqueness assertion, as desired. ∎
By combining Proposition 5.3 and Theorem 6.4, (ii) and (iii), we obtain the following assertion.
Corollary 6.6.
Let us keep the above notation and assumptions, and suppose that $\sum\limits_{j=1}^{n}a_{i}^{[j]}<p^{N}$ for every $i=1,\cdots,r$. Then, the following assertions hold.
  • (i)
    $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$ may be represented by a (possibly empty) stack over $S$ whose structure morphism $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}\rightarrow S$ is (not representable but) quasi-finite and proper. Moreover, if $n=2$, then this morphism is faithfully flat.
  • (ii)
    $\mathcal{U}^{F}_{\mathscr{X}^{(N)},\omega,\mathcal{L}}$ is nonempty if and only if $\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},n,\rho(\vec{a})}$ is nonempty and $n$ divides the integer $\frac{n(n-1)(2g-2+r)}{2}+\sum\limits_{i=1}^{r}\sum\limits_{j=1}^{n}a_{i}^{[j]}$.
Next, we perform an explicit computation of the number of maximally $F^{(1)}$-destabilized parabolic bundles of rank $2$. In the non-parabolic case, this computation was carried out in  [Wak1] and  [Wak3] by applying a correspondence with dormant $\mathrm{PGL}_{2}$-opers.
Theorem 6.7 (cf. Theorem C).
Let us take a line bundle $\mathcal{L}$ on $X^{(1)}$ of even degree and an element $\vec{a}:=(a_{1},\cdots,a_{r})\in(\Xi_{2,1}^{<})^{\times r}$, where $a_{i}:=(a_{i}^{[1]},a_{i}^{[2]})$ ($i=1,\cdots,r$). We write $\omega:=(2,\vec{1}_{\times 2},\vec{a}/p)$, and suppose that $r+\sum\limits_{i=1}^{r}(a_{i}^{[1]}+a_{i}^{[2]})$ is even and $\sum\limits_{i=1}^{r}(a_{i}^{[2]}-a_{i}^{[1]})<2g-2+r<\frac{p}{2}$. Then, the cardinality $\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))$ of the set of isomorphism classes of objects in $\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k)$ satisfies the following inequality:
$\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))\leq 2\cdot p^{g-1}\cdot\sum_{j=1}^{p-1}\frac{\prod\limits_{i=1}^{r}(-1)^{(j+1)(a_{i}^{[2]}-a_{i}^{[1]}+1)}\sin\left(\frac{(a_{i}^{[2]}-a_{i}^{[1]})j\pi}{p}\right)}{\sin^{2g-2+r}\left(\frac{j\pi}{p}\right)}.$ (6.22)
Moreover, this inequality becomes an equality when $\mathscr{X}$ is sufficiently general, or more precisely, classified by a point of ${{}^{\circledcirc}}\mathcal{M}_{g,r}$ (cf. (5.19)) for $(n,N)=(2,1)$.
Proof.
By Corollaries 5.4, 6.6, and  [Wak3, Theorem 7.41], we have
$\sharp(\mathcal{U}^{F}_{\mathscr{X}^{(1)},\omega,\mathcal{L}}(k))\leq 2\cdot p^{g-1}\cdot\sum_{j=1}^{p-1}\frac{\prod_{i=1}^{r}\sin\left(\frac{(2\cdot\tau(a_{i}^{[2]}-a_{i}^{[1]})+1)j\pi}{p}\right)}{\sin^{2g-2+r}\left(\frac{j\pi}{p}\right)},$ (6.23)
where, for an integer $b$ with $0\leq b\leq p-1$, we set $\tau(b):=\frac{b-1}{2}$ (resp., $\tau(b):=\frac{p-1-b}{2}$) if $b$ is odd (resp., even). On the other hand, the following equality holds:
$\sin\left(\frac{(2\cdot\tau(a_{i}^{[2]}-a_{i}^{[1]})+1)j\pi}{p}\right)=(-1)^{(j+1)(a_{i}^{[2]}-a_{i}^{[1]}+1)}\cdot\sin\left(\frac{(a_{i}^{[2]}-a_{i}^{[1]})j\pi}{p}\right).$ (6.24)
Thus, the first assertion follows from (6.23) and (6.24).
The second assertion follows from the generic étaleness of $\overline{\mathcal{O}}p^{{}^{\mathrm{Zzz...}}}_{\mathscr{X},2,\rho(\vec{a})}/S$ proved in  [Moc, Chap. II, Theorem 2.8]. ∎
Acknowledgements
We are grateful for the many constructive conversations we had with algebraic curves in positive characteristic, who live in the world of mathematics! Our work was partially supported by Grant-in-Aid for Scientific Research (KAKENHI No. 21K13770).